### v1

# ESGSenticNet: A Neurosymbolic Knowledge Base for Corporate Sustainability Analysis

Abstract
Evaluating corporate sustainability performance is essential to drive sustainable business practices, amid the need for a more sustainable economy. However, this is hindered by the complexity and volume of corporate sustainability data (i.e. sustainability disclosures), not least by the effectiveness of the NLP tools used to analyse them. To this end, we identify three primary challenges – immateriality, complexity, and subjectivity, that exacerbate the difficulty of extracting insights from sustainability disclosures. To address these issues, we introduce ESGSenticNet, a publicly available knowledge base for sustainability analysis. ESGSenticNet is constructed from a neurosymbolic framework that integrates specialised concept parsing, GPT-4o inference, and semi-supervised label propagation, together with a hierarchical taxonomy. This approach culminates in a structured knowledge base of 44k knowledge triplets – (‘halve carbon emission’, supports, ‘emissions control’), for effective sustainability analysis. Experiments indicate that ESGSenticNet, when deployed as a lexical method, more effectively captures relevant and actionable sustainability information from sustainability disclosures compared to state of the art baselines. Besides capturing a high number of unique ESG topic terms, ESGSenticNet outperforms baselines on the ESG relatedness and ESG action orientation of these terms by 26% and 31% respectively. These metrics describe the extent to which topic terms are related to ESG, and depict an action toward ESG. Moreover, when deployed as a lexical method, ESGSenticNet does not require any training, possessing a key advantage in its simplicity for non-technical stakeholders.
keywords:
Generative AI, Large Language Models, Knowledge Base, Lexicons, Corporate Sustainability Analysis
journal: Information Processing and Management
[…]
[1]organization=College of Design and Engineering, National University of Singapore,addressline=9 Engineering Drive 1, postcode=117575, country=Singapore
[…]
[2]organization=School of Computing, National University of Singapore,addressline=13 Computing Drive, postcode=117417, country=Singapore
[…]
[3]organization=Asian Institute of Digital Finance, National University of Singapore,addressline=Innovation 4.0, 3 Research Link, #04-03, postcode=117602, country=Singapore
[…]
[4]organization=College of Computing and Data Science, Nanyang Technological University,addressline=50 Nanyang Ave, postcode=639798, country=Singapore
[…]
[5]organization=Institute of High Performance Computing, Agency for Science, Technology and Research,addressline=Fusionopolis Way, #16-16 Connexis, postcode=138632, country=Singapore
[…]
[6]organization=Sustainable and Green Finance Institute, National University of Singapore,addressline=Innovation 4.0, 3 Research Link, #02-02, postcode=117602, country=Singapore
[…]
[7]organization=NUS Business School, National University of Singapore,addressline=Mochtar Riady Building, 15 Kent Ridge Dr, Mocthar Riady Building, postcode=119245, country=Singapore
[…]
[8]organization=Honorary Research Fellow, Imperial College London,
1 Introduction
With growing concerns over social inequality and global warming, it is imperative to ensure the sustainability of public companies [13]. To this end, accurate insights into the sustainability efforts of corporations are crucial to drive responsible business practices [40]. However, the complexity and volume of corporate sustainability data [38], often disclosed through sustainability reports (i.e. sustainability disclosures), can make these insights inaccessible without specialised tools.
While NLP tools can unlock insights from this vast array of corporate sustainability data, their deployment remains nascent within this field, often lacking the specialised design needed for corporate sustainability analysis. Specifically, they do not address the challenges of immateriality, complexity and subjectivity, which are frequently associated with sustainability data [41].
  • Immateriality: Sustainability disclosures are lengthy, containing significant amounts of irrelevant content for sustainability analysis [51, 28]. Key or material information for stakeholders can often be obfuscated [49], making their extraction via NLP tools challenging.
  • Complexity: Sustainability ideas are sophisticated, and are typically expressed by complex syntactic constructions [50]. The distinctiveness in sustainability language necessitates that NLP tools must encompass more than just general language understanding capabilities.
  • Subjectivity: Stakeholders are interested in diverse sustainability areas, and utilise different frameworks and methods for sustainability analysis [39, 46]. Therefore, designing an NLP tool that comprehensively meets the needs of different stakeholders is challenging.
As a result, NLP algorithms, including the state of the art, are largely ineffective at extracting useful corporate sustainability insights. In our experiments later in this paper (section 6.3), we validate this ineffectiveness by benchmarking these algorithms in the sustainability topic analysis task.
In response to these limitations, we develop ESGSenticNet, a publicly available knowledge base that comprises 44k sustainability knowledge triplets from 23k unique concepts, specifically designed to mine insights from sustainability disclosures. To address the aforementioned issues of complexity, immateriality and subjectivity, ESGSenticNet is constructed from a neurosymbolic framework that blends symbolic linguistics rules with sub-symbolic reasoning through deep learning. The symbolic component of this framework involves specialised concept parsing that leverages predefined linguistic patterns and rules within the sustainability domain to identify meaningful concepts within sustainability text, addressing the complexity of sustainability language. Additionally, these rules focus on the most pertinent linguistic structures in sustainability, enabling the identification of key sustainability information such as sustainability efforts and actions to reduce immateriality [7]. Building on the structured outputs of the symbolic component, the subsymbolic component integrates GPT-4o inference with semi-supervised label propagation to classify the relation between the parsed concepts and their respective sustainability categories. By interpreting nuanced language variations and contextual subtleties that symbolic rules cannot fully capture, this component ensures that concepts are organised according to key sustainability categories, further enhancing their materiality. Additionally, these categories within ESGSenticNet are arranged according to a hierarchical and comprehensive taxonomy, addressing subjectivity by allowing concepts to be leveraged at different levels of granularity, tailored to the specific interest of stakeholders. The result is an extensive knowledge base, ESGSenticNet, that comprises actionable information in the form of knowledge triplets. These triplets highlight concepts and their relation toward sustainability themes, such as (‘minimise paper waste’ (concept), supports (relation), ‘waste management’ (category)).
We evaluate ESGSenticNet’s ability to extract meaningful insights from sustainability disclosures, by deploying the knowledge base as a structured lexicon method for sustainability topic analysis. In this task, the concepts within ESGSenticNet’s knowledge triplets are matched from a large sustainability corpora. ESGSenticNet improves on established baselines significantly, more extensively capturing sustainability related information, including material content that highlight sustainability efforts. ESGSenticNet yields a significant number of ESG topic terms, outperforming existing state of the art baselines by at least 26% and 31% respectively on the ESG relatedness and ESG action orientation of these terms. This highlights demonstrates ESGSenticNet’s potential to contribute meaningfully to sustainability analysis. Moreover, as a lexical method, ESGSenticNet is able to extract ESG topic terms without any training or significant technical expertise. This significantly allows it to be leveraged by stakeholders of different backgrounds.
The rest of our paper is arranged as follows. We first develop a comprehensive and hierarchical sustainability taxonomy for organising our concepts (section 4). Secondly, we delve into the extraction of concepts via ESG Concept Parser (section 5.2), processing the concepts to ensure coherence (section 5.3), as well as our methodology for classifying these concepts via GPT-4o and semi-supervised label propagation (sections 5.45.55.65.7). Finally, we evaluate the accuracy of ESGSenticNet labels (section 5.9), before running extensive experiments to assess the effectiveness of ESGSenticNet for sustainability analysis against established baselines (section 6). The development process for ESGSenticNet is summarised by figure 1, showcasing the key steps for concept parsing and taxonomy derivation, and how they are intertwined with the labelling process (1-4).
2 Literature Review
NLP-driven Sustainability Analysis. To enhance sustainability analysis, NLP can be deployed for different tasks, such as analysing sustainability topics [65], sustainability sentiment classification [43], extracting textual patterns [8], among other tasks. The insights extracted from NLP can powerfully reduce time costs and increase the consistency of evaluation [2, 15], making NLP a useful tool for sustainability analysis. Yet, despite NLP’s potential applications, it has only been explored within sustainability analysis to a limited degree, often lacking specialised design [41]. This motivates the development of specially designed NLP tools for corporate sustainability analysis.
Lexicons for Sustainability Analysis. For corporate sustainability analysis, lexical approaches have been widely leveraged for textual analysis [22], due to their simplicity and relevance to financial text analysis [53, 55]. To facilitate lexical analysis, sustainablity-specific lexicons have been developed. Approaches such as Baier’s dictionary [6], have focused on developing single-word lexicons according to ESG pillars. However, while Baier’s dictionary has been a popular choice for sustainability analysis works [25], single word lexicons cannot adequately capture the semantics of the sustainability field. In contrast, dictionaries that have collected sustainability related text spans (of multiple-grams) [17, 29], are limited, not publicly available, and not comprehensively labelled beyond ESG pillars. Additionally, due to their reliance on manual human-annotation, these dictionaries have an exceedingly small number of lexicons (less than 800), raising questions on their comprehensiveness, and consequently their effectiveness for capturing textual patterns. This motivates the development of an extensive knowledge base that can function as a structured dictionary of concepts, and the exploration of automated methods such as NLP for constructing and organising sustainability lexicons.
Knowledge Base Construction. Methods for constructing knowledge bases can vary according to the domain and application [26]. Notable knowledge bases, such as ConceptNet [31] involved concept parsing and human annotation, SenticNet 5 [11] leveraged deep learning and lexical substitution, question answering via advanced graphical neural networks (GNNs) [63]. Newer avenues for knowledge construction involve generative-LLMs, which have shown strong performance in related tasks such as reasoning, construction, and generalisation [67]. Unlike all the previous methods which heavily rely on semantic similarities for labelling, generative LLMs possess powerful emergent capabilities that extend beyond textual similarities [58]. These include reasoning, knowledge relationship construction, and generalisation to different domains [67]. Moreover, unlike GNNs that rely on heavily structed text data [32], generative-LLMs can flexibly adapt to unstructured text. This, coupled with added domain flexibility and inferencing capabilities on limited textual cues, motivates our use of generative-LLMs for knowledge derivation within sustainability.
3 Data
1680 sustainability disclosures from a wide range of Singapore Exchange (SGX) based companies, in the period of 2017 to 2022 inclusive, are utilised for taxonomy and knowledge base construction. These reports can be found online via respective company websites, and on request, we can provide a list of the reports utilised (i.e. the companies and year of reporting).
Figure 1: Overview of ESGSenticNet Development
4 Sustainability Taxonomy
Firms may not report in accordance with existing sustainability taxonomies (i.e. GRI [33]), given that their operations are complex and may not be adequately captured by these taxonomies [41]. To further align taxonomies with sustainability report content, data-driven approaches can be utilised for taxonomy derivation [5]. Yet, empirical approaches in the literature are limited, and these methods do not involve human-in-the-loop inference to ensure the meaningfulness of constructed sustainability topics. These challenges motivate our taxonomy construction, that, while empirically driven, involves human-in-the-loop inference to ensure the relevance and meaningfulness of categories.
4.1 Methodology
Sustainability disclosures are tokenised into separate sentences, with each sentence transformed into embeddings by Sentence-BERT (S-BERT) [45]. BERTopic is then run on these embeddings, with the configuration of KeyBERT [20]. The topic words are then provided to sustainability experts (A) to facilitate their development of the taxonomy for ESGSenticNet (table 1). Examples of topic words and their inferred topics can be found in table 1. The result is a taxonomy (table 2) with the following features.
  • Realistic: Given that taxonomy development leverages empirical insights from the sustainability reports themselves, categories can capture content within sustainability reports that might otherwise be overlooked by standard frameworks. For instance, ‘Green Technology’ is derived as a taxonomy category, while not being explicitly emphasised in standard reporting frameworks (i.e. GRI [33]).
  • Activities Cross-Theme: To capture the interconnectedness of sustainability dimensions [38], categories include areas of a firm’s activities (‘operations’, ‘development’ etc.), on top of sustainability impacts (‘emissions control’, ‘waste management’ etc.). In contrast with existing frameworks, this offers insight into the sustainability of specific business activities.
  • Granularity: The taxonomy has a hierarchical nature to enable analysis at different levels of granularity.
Topic keywords Category
strengths, staff, training, opportunities, employees Workplace Development
cybersecurity, security, cybersecure, authentication Data Privacy & Cybersecurity Protection
sustainability, technology, technological, digitalisation Green Technology
sponsorship, sponsor, charity, scholarship Community Empowerment
Table 1: Examples of Categories Inferred from Topic Keywords
4.2 Resultant Taxonomy
Table 2 shows the hierarchical taxonomy derived for ESGSenticNet. It contains the definitions of the pillars (Environmental, Social, Governance) and topics used in ESGSenticNet. Along with the pillars, the broad topics, and cross-broad topics are bolded, with their constituent sub-topics highlighted in bullet points. Beside the broad and cross-broad topics, a letter denotes which pillar the topic falls under – (E) Environmental, (S) Social, (G) Governance.
Category Description
Environmental Focuses on impact on the environment. This includes, but is not limited to, green innovation, green infrastructure, resource stewardship, energy consumption, waste management, pollution reduction, natural resource conservation, and animal welfare.
Social Focuses on managing relationships with its employees, suppliers, customers, and the communities. This includes, but is not limited to, employee diversity, equity, working conditions, customer relations, partnerships and social engagement.
Governance Focuses on the system of rules, practices, and processes by which a company is directed and controlled. This includes, but is not limited to, corporate oversight, corporate transparency, executive compensation, ethical behavior, and shareholder rights.
Development (E) Drives innovation and growth in sustainability by creating new technologies or infrastructure for sustainability, or supporting these efforts through green financing.
$\bullet$ Green Technologies Technological systems, products, and processes that assist in reducing environmental impact.
$\bullet$ Green Infrastructure Environmentally sustainable buildings, infrastructure, energy solutions and urban systems.
$\bullet$ Green Financing Financial services that support efforts aimed at environmental sustainability.
Operations (E) Focuses on optimizing existing business activities for the sole purpose of minimizing environmental impact. This includes minimising environmental impact by enhancing current practices in corporate operations, supply chains, production processes, or land management, distinct from the development of innovative solutions and technologies.
$\bullet$ Sustainable Corporate Operations Environmentally responsible practices within corporate offices. Examples include but are not limited to promoting sustainability amongst employees, eco-friendly office lighting.
$\bullet$ Sustainable Supply Chain Logistics and distribution methods or practices that are environmentally responsible, ensuring sustainability in the movement and handling of goods.
$\bullet$ Sustainable Production Processes Manufacturing processes or practices that are environmentally responsible.
$\bullet$ Sustainable Land Management Using and managing land in a way that is environmentally responsible, conserves natural resources and protects biodiversity.
Ecological Conservation (E) Manages the company’s interaction with natural ecosystems, focusing on minimizing harmful impacts and promoting biodiversity and habitat conservation.
Emissions Control (E) Targets the reduction of various emissions from company operations, with an emphasis on mitigating climate change impacts or protecting air quality.
$\bullet$ Climate Emissions Manages greenhouse gas emissions that contribute to global warming, such as CO2 and methane. The phrase must explicitly relate to greenhouse gases or related terms such as ’carbon’, ’methane’.
$\bullet$ Air Quality Emissions Manages emissions of non-greenhouse gases, like toxic gas, particulates and sulfur dioxide, which are critical for local air quality and public health.
Waste Management (E) Concentrates on managing waste after it has been produced, including the reduction, handling, disposal, and treatment of solid waste and wastewater. It involves but is not limited to recycling, reusing materials, and employing wastewater treatment processes to mitigate environmental degradation and facilitate resource recovery.
$\bullet$ Wastewater Management Focuses on the treatment, reuse and recycling of wastewater or liquid waste to prevent pollution or conserve water resources.
$\bullet$ Solid Waste Management Addresses the management of solid wastes, promoting reduction, recycling treatment, and responsible disposal to minimize environmental impact and encourage circular economy principles.
Resource Optimisation (E) Prioritizes the proactive, efficient, and sustainable selection and use of energy, water, and materials. This approach focuses on minimizing environmental impact through the strategic choice and utilization of resources. Distinct from waste management as it does not primarily deal with waste but rather prevents waste generation by optimizing resource use from the start.
$\bullet$ Renewable and Efficient Energy Focuses on using energy efficiently, reducing energy consumption and increasing the use of renewable energy sources such as solar, wind, and hydroelectric power to minimize environmental impact.
$\bullet$ Material Sustainability Prioritizes the use of sustainable and renewable materials, focusing on maximizing their efficiency from the outset. Distinct from recycling and reuse, which are part of waste management, this approach aims to reduce environmental impact before materials enter the waste stream.
$\bullet$ Water Conservation Focuses on using water efficienctly, reducing water usage and enhancing sustainable water management practices to minimize initial consumption. Distinct from water recycling and reusing, which address management of wastewater.
Workplace (S) Focuses on the social aspects of sustainability within the company to promote equity, inclusivity, or quality of life, emphasizing employee well-being, inclusivity, or professional growth.
$\bullet$ Workplace Wellness Efforts that focus on the satisfaction, emotional and physical well-being of employees.
$\bullet$ Workplace Diversity Efforts that emphasise gender and racial diversity and fairness in the workplace, such as anti-discrimination measures, fair hiring practices etc.
$\bullet$ Workplace Development Efforts that focus on the skill training and professional development of staff.
Communications (E/S/G) Focuses on how a company transparently and honestly communicates and discloses its activities.
Outreach (S) Enhances social sustainability by contributing to community well-being, fostering customer satisfaction, building customer relationships, or collaborating with strategic partners to promote equity, inclusivity, and quality of life.
$\bullet$ Community Empowerment Contributions to the welfare of local communities through various forms of support and engagement.
$\bullet$ Strategic Partnerships Collaborations with stakeholders like NGOs, businesses, and charities to achieve shared sustainability goals.
$\bullet$ Customer Engagement Fostering customer satisfaction, strong relationships and reputation with customers.
Management (G) Ensures strategic, ethical, or effective governance through rigorous oversight, diverse board composition, active stakeholder engagement, or high ethical standards within company leadership.
$\bullet$ Management Composition Structures the board or management with diverse expertise and backgrounds to strengthen decision-making and governance effectiveness.
$\bullet$ Management Ethics Enforces strict ethical standards within company leadership, ensuring management actions uphold corporate integrity.
$\bullet$ Management Compensation Aligns management, executive and board remuneration with company performance, sustainability or ethical objectives.
$\bullet$ Stakeholder Engagement Engages key stakeholders to align board or management decisions with broader interests and feedback.
$\bullet$ Oversight Involves the systematic review and evaluation of management actions to ensure they align with the company’s performance goals, sustainability goals and ethical standards.
Compliance (G) Ensures adherence to laws, regulations, or standards across environmental protection, labor rights, consumer safety, ethical conduct, financial practices, or data protection, distinguishing itself from voluntary sustainability initiatives by focusing on mandatory requirements.
$\bullet$ Environmental Compliance Adherence to environmental laws such as for emissions, water pollution, protection of biodiversity etc.
$\bullet$ Worker & Consumer Safety Adherence to laws that guarantee the safety of products and services to consumer health, as well as fair labor practices and the protection of workers’ rights.
$\bullet$ Business Compliance Adherence to anti-corruption, anti-competitive practices, financial regulations, or financial reporting.
$\bullet$ Data Privacy & Cybersecurity Protection Protection and confidentiality of data, which includes measures against cyber threats and compliance with data security regulations.
Table 2: ESGSenticNet Sustainability Taxonomy
5 ESGSenticNet
ESGSenticNet comprises knowledge triplets with the format of (concept, relation, category), following from [27]. To form the knowledge triplets, concepts within our knowledge base are labelled according to their relations (table 3) with the categories in our taxonomy (table 2). The relations, ‘supports’ and ‘undermines’, clarify how concepts advance or impede crucial aspects of corporate sustainability. In contrast, the relation ‘aligns with’ does not consider a concept’s impact, highlighting a concept’s general connection with broader pillars of environmental, social or governance. Additionally, categories are divided according to their different category types–pillar, broad, cross-broad, sub, cross-sub, as shown in table 4, with relations assigned according to rules (section 5.1).
ESGSenticNet is constructed through a pipeline that involves three main phases: parsing concepts from the text corpus (section 5.2), processing parsed concepts to ensure coherence (section  5.3), labelling processed concepts according to their relation with respect to our developed taxonomy (sections 5.55.65.45.7). Concept parsing involves extracting candidate concepts from sustainability disclosures through dependency parsing and part-of-speech tagging. From these candidates, coherent concepts are filtered through text parsing followed by leveraging GPT-4 [3] (section 5.2). The filtered concepts undergo a labelling methodology that we have developed to minimise computational cost while enhancing the diversity of labelled concepts. Accordingly, this labelling methodology (figure 1 1 through 4) entails leveraging semantic graphs, clustering techniques, seed selection algorithms, and label propagation. This methodology allows us to label a large amount of diverse concepts despite performing inference on a subset of the concept samples. As a result, the extensiveness and diversity of our knowledge base is significantly increased, which is imperative to its application for corporate sustainability analysis.
5.1 ESGSenticNet Relations & Rules
Relation Description
aligns with The concept is positively or negatively related to a specific Environmental, Social, or Governance (ESG) pillar.
supports The concept directly advances the sustainability topic with a clear and immediate positive impact.
undermines The concept directly impedes the sustainability topic with a clear and immediate negative impact.
Table 3: Relations Types
Category Type Categories
Pillar ‘Environmental’, ‘Social’, ‘Governance’.
Broad ‘Development’, ‘Operations’, ‘Workplace’, ‘Outreach’, ‘Management’, ‘Communications’, ‘Compliance’
Sub Children categories of broad categories
Cross-Broad ‘Resource Optimisation’, ‘Waste Management’, ‘Emissions Control’, ‘Ecological Conservation’
Cross-Sub Children categories of cross-broad categories
Table 4: Categories and category types
ESGSenticNet is defined with the following relations (table 3), category types (table 4), as well as the following rules.
  • Pillar Assignment: Each concept ‘aligns with’ a specific pillar, mapping the concepts to the broader sustainability context. Single Label within Category Types: Each concept can hold only one relation with one category within each category type.
  • Cross-Labels between Parent & Children: Each concept can hold relations with higher-level categories (‘broad’ and ‘cross-broad’), as well as with more detailed categories (‘sub’ and ‘cross-sub’).
  • Cross-Labels between Cross & Non-Cross: Concepts that ‘align with’ the ‘environmental’ pillar can hold relations with ‘cross-broad’ and ‘cross-sub’ categories, in addition to ‘broad’ and ‘sub’ categories. This reflects the multi-faceted nature of environmental concepts, which can simultaneously impact environmental stewardship (represented by ‘cross-broad’, ‘cross-sub’ categories) and strategic business activities (represented by ‘broad’ and ‘sub’ categories).
5.2 ESG Concept Parser
Multiword expressions are linguistic phenomena that refer to common linguistic constructions such as idioms, collocations and sayings [16]. In the context of sustainability analysis, multiword expressions frequently communicate key sustainability ideas [51]. We refer to these multiword expressions as sustainability ”concepts”. Extracting these concepts or ”concept extraction” has been explored for metaphors [36], finance [18] and sentiment analysis [10]. Yet, to the best of our knowledge, concept extraction within the sustainability domain has not been studied. Therefore, to capture these sustainability concepts, a new concept parser is required, with an explicit focus on the distinct elements of sustainability discourse. These include the following designs:
  • Complex Nominals, Modifiers, Nouns & Adjectives: Sustainability discourse is complex, containing a high concentration of adjectives and nouns which commonly function as modifiers within nominal phrases [50]. Phrases such as ‘renewable energy’, ‘carbon emissions’, demonstrate constructions that efficiently convey complex sustainability ideas.
  • Descriptive & Prescriptive Elements: Sustainability discourse often involves descriptive and prescriptive language. The former (i.e. stative verbs like ‘invested’, ‘reduced’) is utilised when a company reports its current sustainability activities [42], while the latter (i.e. verbs of volition like ‘plan’, ‘commit’) is observed when a company pledges future sustainability activities [48].
To effectively extract these unique aspects of sustainability language, we formalise sustainability concepts, or ESG concepts, through the following syntactic structure–verbs acting on complex noun phrases, where the words involved function as modifiers, descriptive and prescriptive elements. Specifically, the concepts are two or three word phrases, typically beginning with a $verb$, followed by $noun+noun$, or $adjective+noun$ combinations with specific grammatical dependencies. These dependencies include: $verb-noun$ connection via (nsubj, obj, obl), $adjective-noun$ or $noun-noun$ connection via (compound, amod, nn, appos, flat, nmod). This framework underpins the ESG Concept Parser algorithm (1), designed to capture key ideas within sustainability disclosures, particularly sustainability efforts. For instance, ‘improve (VB) workplace (NN) safety (NN)’, where ‘workplace’ is the $compound$ of ‘safety’ and ‘safety’ is the $dobj$ of ‘improve’.
Algorithm 1 ESG Concept Parser
1:Input: Sentence from sustainability disclosure
2:Output: ESG Concepts
3:POS tag the sentence
4:Dependency parse the sentence
5:for dependency, head, dependent in sentence do
6:    if dependency $\in\{\textit{nsubj},\textit{obj},\textit{obl}\}$ then
7:         if head is noun & dependent is verb then
8:             add to Phrases: dependent+head as verb+noun combination
9:         end if
10:    end if
11:end for
12:for $word_{1}$, $word_{2}$ in Phrases do
13:    for dependency, head, dependent in sentence do
14:         if head is noun & dependency $\in\{\textit{compound},\textit{amod},\textit{nn},\textit{appos},\textit{flat},\textit{nmod}\}$ then
15:             if dependent is noun then
16:                 add to Phrases: $word_{1}$+dependent+$word_{2}$ as verb+noun+noun combination
17:             end if
18:             if dependent is adj then
19:                 add to Phrases: $word_{1}$+dependent+$word_{2}$ as verb+adj+noun combination
20:             end if
21:         end if
22:    end for
23:end for
5.3 Processing concepts for coherence
To minimise computational cost, parsed ESG concepts from algorithm (1) are filtered to retain the top 110k most frequently occurring ones. Of these, we eventually annotate 23k concepts based on our estimated computational resources, which dictates the number of seeds we select within our semantic graph. Prior to the annotation process, these concepts must undergo processing to ensure their coherence. This involves utilising GPT-4o in few-shot setting (Brown, 2020) to ensure their coherence. GPT-4o is leveraged due to the capacity of large generative LLMs to infer [58], as well as their powerful natural language understanding capabilities [62]. Using GPT-4o to process concepts for coherence involves conducting the following inference tasks in order: (1) reorder words, if required, to improve the coherence of the ESG concept, (2) determine if the subsequent ESG concept is intelligible. We provide the prompt for (1) and (2) through figures 3 and 4 respectively.
[Task Description] As a language expert, your task is to rearrange words in phrases to enhance their coherence and intelligibility if necessary. [Response Instructions] Please list all your outputs for all the phrases given. Follow these steps for each given phrase: Assess whether reordering the words in the phrase can improve its coherence and intelligibility. If needed, reorder the words to achieve the best clarity and coherence, but do not add new words and do not change the words. If no reordering is needed, retain the original order. Record the output accordingly. [Response Format] Fill in your response in the triple backticks below using the following format: Input: <original phrase> Output: <reordered or original phrase>
Figure 2: Prompt for reordering an ESG concept for coherence
[Task Description] As a language expert, you are tasked with determining if a phrase is intelligible. [Response Instructions] Please list all your outputs for all the phrases given. Follow these steps for each given phrase: Assess if a given phrase is intelligible. If the phrase is intelligible, output a ‘True’ for the phrase. If the phrase is not intelligible, output a ‘False’ for the phrase. [Response Format] Fill in your response in the triple backticks below using the following format: Input: <phrase> Output: <True/False>
Figure 3: Prompt for evaluating an ESG concept’s coherence
5.4 Semantic Graph Construction
$G=(V,E)$ (1)
$V=\{\mathbf{e}_{i}\mid\mathbf{e}_{i}=\text{S-BERT}(cp_{i})\}$ (2)
$\text{sim}(\mathbf{e}_{i},\mathbf{e}_{j})=\frac{\mathbf{e}_{i}\cdot\mathbf{e}_{j}}{\|\mathbf{e}_{i}\|\|\mathbf{e}_{j}\|}$ (3)
$E=\{(\mathbf{e}_{i},\mathbf{e}_{j})\mid\text{sim}(\mathbf{e}_{i},\mathbf{e}_{j})>0.80\}$ (4)
Processed ESG concepts undergo labelling to derive their relationships with different sustainability topics. Given the exceedingly large number of ESG concepts, annotating all concepts would be prohibitively expensive. To overcome this, we devise a semi-supervised framework that enables a large volume of diverse ESG concepts to be labelled from annotating only a subset of all ESG concepts. Via a semantic graph, we leverage the semantic similarities between ESG concepts to propagate labels from annotated concepts (seeds) to unlabelled concepts (non-seeds). This graph is constructed by transforming each concept $cp_{i}$ into embeddings $e_{i}$ through S-BERT [45], with edges constructed by leveraging a cosine similarity threshold of 0.80 between each node $e_{i}$.
5.5 Seed Selection
Within our semantic graph, seeds are selected for annotation through data clustering and graph analysis. Given the semi-supervised label propagation from seeds to non-seeds, we formulate the following objectives for seed selection: (I) Seeds must be semantically diverse to ensure that diverse labelled samples are yielded post label propagation. This enhances the variety and comprehensiveness of labelled ESG concepts in our knowledge base. (II) Seeds must propagate their labels to as many non-seeds as possible. This maximises the number of labelled samples we can attain while minimising annotation costs.
5.5.1 Data Clustering
$CQI=\frac{|\{e\in V\mid\alpha_{e}>0.60\}|}{|V|}$ (5)
To accomplish (I), the embeddings of ESG concepts $e_{i}$ are clustered to enable seed selection from diverse semantic spaces. This clustering involves projecting the embeddings to a lower dimensional space via UMAP [37], before running HDBSCAN [12] on these projected embeddings. While HDBSCAN is flexible to the shape and number of clusters, and robust to noise [12], reducing the dimensionality of embeddings ensures good clustering performance. This is because HDBSCAN relies on distance-based metrics which lack meaningfulness in the high dimensional spaces of textual embeddings [4]. The embeddings are projected to lower dimensions for the sole purpose of clustering. In our later section, label propagation leverages the default dimensions of S-BERT embeddings.
Clusters are evaluated via a Confidence Quality Index ($CQI$), which denotes the proportion of embeddings, $e$, out of all embeddings, $V$, with a clustering confidence level $\alpha_{e}$ beyond 0.60. For our use of HDBSCAN, this simple metric is more appropriate than others like Slihouette Score [47], which assumes that all embeddings belong to a cluster. Hyperparameter tuning is done via random search, with the best run having a $CQI$ of 0.66, UMAP parameters $n$_$neighbours$$=$32, $n$_$components$$=$2, and the HDBSCAN parameter $min$_$cluster$_$size$$=$2.
5.5.2 Graph Analysis
$\begin{split}Q(\mathbf{e}_{i})=|\left\{(\mathbf{e}_{i},\mathbf{e}_{j})\in E\mid(\mathbf{e}_{i}\in\mathbf{c_{k}})\right.\\ \left.\wedge\left(\forall\mathbf{e}_{s}\in S_{c_{k}},(\mathbf{e}_{s},\mathbf{e}_{j})\notin E\right)\right\}|\end{split}$ (6)
To accomplish (II), graph analysis is utilised to select seeds that are the most consequential nodes. Given that edges facilitate label propagation, a node is considered more consequential if it possesses a greater number of unique edges. Unique edges are defined as connections to nodes that do not share an edge with the already selected seed nodes within the same cluster. We formulate the local consequential score $Q(e_{i})$, where $S_{c_{k}}$ is the set of nodes already selected as seeds within the same cluster, $c_{k}$, as the node $e_{i}$.
5.5.3 Seed Selection Algorithm
Capitalising on Data Clustering and Graph Analysis, an algorithm (2) is developed for seed selection to achieve objectives (I) & (II). Given a target number of seeds to label, $T$, and the embeddings of all ESG concepts, $V$, seeds, $S_{total}$, are selected from clusters, $C$. Each cluster, $c_{k}$, within $C$, is a set of embeddings of concepts, $e_{i}$. Within each $c_{k}$, seeds of a cluster, $S_{c_{k}}$, are selected by taking the embedding with the maximum local consequential score $Q(e_{i})$ at each selection step. This maximises the selection of seeds with unique edges to increase label propagation to non-seeds. Additionally, each cluster should ideally possess an equivalent proportion, $P$, of seeds, and at least one seed each. We accomplish this by iteratively adjusting $P$ according to parameter $\beta$ (ESGSenticNet uses $\beta=0.01$), to ensure that our estimated number of seeds $N_{s}$ is as close to our target number as possible $T$ (the difference between $N_{s}$ and $T$ denoted by $\Delta$). From $P$, the specific number of seeds for each cluster, $n_{c_{k}}$ can be computed. This ensures that seeds are fairly represented across different clusters to enable the semantic diversity of seeds.
Algorithm 2 Seed Selection
1:Input: $V$, $T$, $C$
2:Output: $S_{\text{total}}$
3:$N\leftarrow$ $|V|$
4:$P\leftarrow\frac{T}{N}$
5:$\Delta_{prev}\leftarrow\infty$
6:while True do
7:    $N_{s}\leftarrow\sum_{c\in C}\max(1,\lfloor P\cdot|c|\rfloor)$
8:    $\Delta_{curr}\leftarrow|N_{s}-T|$
9:    if $\Delta_{curr}\geq\Delta_{prev}$ then
10:         break
11:    else
12:         $\Delta_{prev}\leftarrow\Delta_{curr}$
13:    end if
14:    if $N_{s}<T$ then
15:         $P\leftarrow P+\beta$
16:    else
17:         $P\leftarrow P-\beta$
18:    end if
19:end while
20:for $c_{k}\in C$ do
21:    $n_{c_{k}}\leftarrow\max(1,\lfloor P\cdot|c_{k}|\rfloor)$
22:    $S_{c_{k}}\leftarrow\emptyset$
23:    while $|S_{c_{k}}|<n_{c_{i}}$ do
24:         $e_{\max}\leftarrow\arg\max_{e\in c_{k}}Q(e)$
25:         $c_{k}\leftarrow c_{k}\setminus\{e_{\max}\}$
26:         $S_{c_{k}}\leftarrow S_{c_{k}}\cup\{e_{\max}\}$
27:    end while
28:    $S_{total}\leftarrow S_{total}\cup S_{c_{k}}$
29:end for
30:return $S_{\text{total}}$
5.5.4 Effectiveness of ESGSenticNet Seed Selection Algorithm
Seed Selection Method Propagated Labels
Highest Degree Centrality 4686
Random Selection 7456
ESGSenticNet 9617
Table 5: Effectiveness of Label Propagation, based on the number of newly propagated labels
In this baseline study, we assess the extent of label propagation through ESGSenticNet’s seed selection algorithm, against naive methods like selecting seeds randomly or based on highest degree centrality scores. Each method is run on the constructed semantic graph (section 5.4), to select different batches of 10k seeds. Aligning with majority of the classification tasks in the main study, three different psuedo-labels are used, and randomly assigned to the selected seeds. Label propagation is run to observe the number of newly propagated labels. Although propagation would be more accurate with real labels, random pseudo-labelling provides an unbiased estimate of propagation extensiveness irrespective of label content, and reduces the cost of extensive testing. Moreover, the entire process is repeated over 100 trials, showing that across each method, the number of newly propagated labels remain consistent despite the randomness of seed labelling. Compared to naive methods, we observe that ESGSenticNet’s seed selection algorithm results in a greater number of newly propagated labels.
5.6 Seed Annotation
Figure 4: Flowchart depicting inference process
Through GPT-4o few-shot prompting, selected seeds are annotated with their relations (‘aligns with’, ‘supports’, ‘undermines’) with respect to a pillar or category, before their labels are propagated within the semantic graph. Similar to popular sustainability lexicons [6, 54], ESGSenticNet’s approach focuses on seed annotation before leveraging textual similarities to expand the initial seed corpus. Yet, while approaches in the literature annotate a small number of seeds, ESGSenticNet annotates an extensive number of seeds ($>16k$) through GPT-4o. Consequently, ESGSenticNet retains a significant proportion of seed concepts directly labelled through GPT-4o, while non-seeds that rely on textual similarity for labelling are less prevalent. This enriches the diversity of ESGSenticNet, given that the reasoning capabilities of GPT-4o [24] allows it to infer relations for concepts outside of similar semantic spaces. For example, consider the concepts ‘use led light’, ‘achieve energy efficiency’ categorised under ‘Resource Optimisation’. Textual similarity methods leveraging these for comparison might not categorise ‘emphasise water importance’ similarly, despite its relevance. However, using GPT-4o, ‘emphasise water importance’ is appropriately labelled, due to the LLM’s capacity to infer semantic relationships beyond mere textual similarities. Moreover, GPT4o as an automated technique allows for a large number of concepts to be labelled, overcoming prohibitive human labelling costs. This allows us to derive an extensive knowledge base.
To improve GPT-4o classification accuracy, inference tasks are divided into smaller, more manageable sub-tasks [14]. These tasks align with the taxonomy’s hierarchy levels, following the flowchart in figure 4. Subsequent categorisations of seeds are contingent upon their initial classification into broader categories, and it aligns with the predefined associations between the ESG pillars and broad categories, as well as the connections between the broad and sub categories. Prompts for these tasks are highlighted in figure 5 (pillar labelling), figure 6 (relation-topic labelling).
[Task Description]
As a sustainability expert, your task is to classify phrases into the categories of Environmental, Social, Governance, or Others based on their explicit relevance to the Environmental, Social, Governance topics.
[Definitions of Categories]
Definitions of ESG according to taxonomy, with the inclusion of ‘Others’
Others: Reserved for phrases that do not meet the explicitness and specificity requirements for Environmental, Social, or Governance. This includes phrases that are vague, ambiguous, or cover multiple Environmental, Social, Governance topics without clear categorization.
[Response Instructions]
Evaluate each phrase’s explicit content to determine its most appropriate category. Assign each phrase to only one category: Environmental, Social, Governance, or Others. Only phrases that are clear and explicit can be categorized under Environmental, Social, or Governance. Phrases that are vague or not clearly related to these topics should be categorized as Others. Ensure all responses strictly adhere to the definitions and format provided.
[Response Format]
Fill in your response in the triple backticks below using the following format:
Input: <phrase>
Response: <Environmental, Social, Governance, or Others>
Figure 5: Prompt for labelling an ESG concept’s pillar
[Task Description] As a sustainability expert, you are tasked with evaluating the relationship between phrases and sustainability-related topics. [Definitions of Relationships] supports: The phrase, specific and explicit to the topic, directly advances it with a clear and immediate positive impact. undermines: The phrase, specific and explicit to the topic, directly impedes its goals with a clear and immediate negative impact. [Definitions of Topics] Definitions of topics according to taxonomy [Response Instructions] Evaluate the explicit content of each phrase to determine the most appropriate and accurate relationship and topic. Identify the single most relevant and appropriate relationship and topic. If no specific and explicit relationship can be determined between the phrase and any topic, or if the phrase does not fit any topic due to vagueness or irrelevance, classify the phrase as ‘not applicable’. [Response Format] For each phrase, provide your analysis using the format below. Output only one tuple if applicable, otherwise state ‘not applicable’. Input: <phrase> Response: <(relationship, topic)> or ‘not applicable’
Figure 6: Prompt for labeling an ESG concept’s relation with a topic
5.7 Label Propagation Algorithm
$L^{(k+1)}=D^{-1/2}AD^{-1/2}L^{(k)},$ (7)
$l_{i}^{(k+1)}=l_{i}^{(0)}\quad\forall i\leq m$ (8)
Annotated seeds have their labels propagated to non-seeds via the label propagation algorithm [56]. This increases the labelled data within our knowledge base without incurring additional labelling costs. Specifically, the following algorithm is iterated until convergence, where $A$ is an adjacency matrix containing $a_{i,j}$ that denotes the edge weight between connected nodes $e_{i}$ and $e_{j}$. Each node represents an embedding of a concept, while the edge weight denotes cosine similarity between concepts. $L$ is a matrix that represents the relations labels of each node with respect to a topic (i.e. supports within the (‘reduce water consumption’, supports, ‘resource optimisation’). $m$ is the number of nodes possessing labels and $D$ is a diagonal matrix containing $d_{i,i}$, which equals the sum of elements in row $i$ of $A$. We use parameters $n\_layers$=50, $\alpha$=0.5.
5.8 ESGSenticNet Statistics
Categories Total Triplets Number of triplets with the relations
supports undermines aligns with
Environmental 5904 - - 5904
Social 6897 - - 6897
Governance 10444 - - 10444
Development (E) 314 298 16
$\bullet$ Green Technologies 58 49 9 -
$\bullet$ Green Infrastructure 105 104 1 -
$\bullet$ Green Financing 16 16 - -
Operations (E) 1470 1209 261 -
$\bullet$ Sustainable Corporate Operations 215 186 29 -
$\bullet$ Sustainable Supply Chain 48 46 2 -
$\bullet$ Sustainable Production Processes 265 148 117 -
$\bullet$ Sustainable Land Management 174 141 33 -
Resource Optimisation (E) 935 855 80 -
$\bullet$ Renewable and Efficient Energy 348 321 27 -
$\bullet$ Material Sustainability 58 50 8 -
$\bullet$ Water Conservation 107 100 7 -
Waste Management (E) 578 493 83 -
$\bullet$ Wastewater Management 82 67 15 -
$\bullet$ Solid Waste Management 166 156 10 -
Emissions Control (E) 533 407 146 -
$\bullet$ Climate Emissions 173 152 21 -
$\bullet$ Air Quality Emissions 36 19 17
Ecological Conservation (E) 238 205 33 -
Workplace (S) 2077 2006 71 -
$\bullet$ Workplace Wellness 463 432 31 -
$\bullet$ Workplace Diversity 237 222 15 -
$\bullet$ Workplace Development 705 704 1 -
Outreach (S) 1212 1211 1 -
$\bullet$ Community Empowerment 498 498 0 -
$\bullet$ Strategic Partnerships 137 137 0 -
$\bullet$ Customer Engagement 77 77 0 -
Management (G) 2332 2305 27 -
$\bullet$ Management Composition 30 30 - -
$\bullet$ Management Ethics 124 118 6 -
$\bullet$ Management Compensation 19 17 2 -
$\bullet$ Stakeholder Engagement 105 105 - -
$\bullet$ Oversight 532 532 - -
Communications (E/S/G) 1196 1192 4 -
Compliance (G) 3405 3093 312 -
$\bullet$ Environmental Compliance 436 403 33 -
$\bullet$ Worker & Consumer Safety 1154 1040 114 -
$\bullet$ Business Compliance 208 170 38 -
$\bullet$ Data Privacy & Cybersecurity Protection 103 100 3 -
Table 6: ESGSenticNet Statistics
Concept Relation Category
build diverse workplace supports workplace diversity
halve carbon emission supports emissions control
involve workplace injury undermines worker & consumer safety compliance
minimise resource consumption supports resource optimisation
organise charity event supports outreach
produce chemical waste undermines operations
Table 7: Samples of knowledge triplets from ESGSenticNet
Samples of knowledge triplets within ESGSenticNet is provided in table 7. Additionally, a full statistical breakdown of ESGSenticNet’s 44k knowledge triplets is outlined in table 6.
5.9 ESGSenticNet Human Evaluation
Sustainability Corpora Total Terms
ESGSenticNet 23245
Baier 491
Naiara 167
Kang* 792
Table 8: Total number of terms of different sustainability corpora, Baier [6], Naiara [44], Kang [29], *indicates not publicly available.
Relation Type Accuracy
Pillar 95.0%
Broad 89.3%
Cross-Broad 86.1%
Sub 87.8%
Cross-Sub 90.1%
Table 9: Accuracy of ESGSenticNet
ESGSenticNet comprises 44232 triplets containing 23245 unique concepts. Of the 23245 concepts, 16011 are seeds and 7234 are non-seeds. We provide a meta statistic of ESGSenticNet (table 8), a full statistical breakdown (table 6), and samples (table 7). To evaluate our knowledge base, human annotators determine the accuracy of relations between concept and categories. 500 concepts are randomly selected from our knowledge base. A concept may be found in multiple knowledge triplets, possessing multiple relations with different categories. The accuracy of the labels are human evaluated at every GPT-4o decision point as shown in figure 4. Specifically, we evaluate the accuracy of relation labels separately for each category type–pillar, broad, cross-broad, sub, cross-sub, with each of them including only select categories (table 4). A provides more details on our annotators and annotation scheme.
For evaluating ESGSenticNet accuracy in section 5.9, the following instructions were given:
The table below presents various concepts along with their polarities and respective relations toward sustainability categories. Each concept may be associated with multiple relations across different sustainability categories. These categories are classified into five category types: pillar, broad, sub, cross-broad, and cross-sub. If a concept has no relevant association with a category, it is marked as ‘not applicable’. Your task is to evaluate whether each concept is correctly related to its corresponding categories. For each category, determine if the relation is the most appropriate compared to the other categories within the same category type. Mark ‘TRUE’ if the relation and category are correct, and ‘FALSE’ if they are incorrect. Additionally, assess the accuracy of the polarity of each concept, and mark ‘TRUE’ if it is correct and ‘FALSE’ if it is incorrect.
6 Topical Analysis of Sustainability Disclosures through ESGSenticNet
Topic analysis of sustainability disclosures provide stakeholders significant insights about a corporation’s sustainability performance [41]. Distinct from discovering topics without any inclination of what the topics should be, topic analysis within sustainability analysis frequently seeks to derive topics within established sustainability themes and frameworks [41]. To elaborate on this, sustainability stakeholders often have specific themes they are monitoring, such as alignment with UN Sustainability Development Goals [57], future orientation of reporting within ESG pillars [23], performance with respect to economic, environmental, social dimensions [54]. Therefore, the demands of the topic analysis within sustainability requires capturing content related to sustainability and its associated frameworks. ESGSenticNet aims to address the shortcomings with respect to unstructured topic terms, unstructured topics, the limited structured methods, as well as the materiality of information extracted. To elaborate, the hierarchical structure of ESGSenticNet ensures that the topics extracted are directly related to a sustainability taxonomy, while the specialised concept parsing and GPT-4o inference increases the materiality and ESG-relevance of extracted topic terms.
6.1 Critique of current methods
Different methods can be leveraged for sustainability topic analysis, including rule-based methods that leverage existing dictionaries [25], and unsupervised topic modelling that can be adapted to explore specific areas within sustainability frameworks [66, 35]. Yet, while having advantages, these methods may not be tailored for sustainability topic analysis, given the specialised nature of sustainability content and the demands of sustainability stakeholders. We elaborate on this in the following.
  • Unrelated topic terms: Existing topic modelling methods, while excellent at producing topic terms without supervision  [1], cannot be tuned to capture specialised word constructions that typically describe sustainability ideas (i.e. nominals, modifiers as described in 5.2[50]. Moreover, sustainability content is multifaceted and complex, often varying across different sectors and contexts [39]. Therefore, captured topic terms may not be relevant to sustainability, leading to topics irrelevant for sustainability analysis.
  • Unstructured topics: Unsupervised topic modelling may not yield topics that are structured according to sustainability themes, even if sustainability-related topics can be derived from topic words. Therefore, the derived topics may not be meaningful for sustainability stakeholders, who typically rely on structured frameworks with predefined sustainability themes and topics to analyse sustainability content [34, 19].
  • Limited Structured Methods: Although publicly available ESG dictionaries such as Baier [6] introduce a structured method for topic analysis [22], they contain lexicons, which we will demonstrate, that are not predominantly related to sustainability. This makes the alignment between the lexicons and their supposed sustainability categories unclear. Other dictionaries (i.e Naira [44]) are limited ($<200$ terms).
  • Material Information: Although topic terms may highlight the presence of sustainability issues, stakeholders often prioritise content that explicitly convey the sustainability efforts or actions undertaken by corporations [7]. Therefore, greater specificity of topic terms is required to enrich the materiality of topical information in alignment with stakeholder interests. Nevertheless, current methods may fall short in generating specialised topic terms that articulate sustainability action, a limitation that we will explore further.
6.2 Topic Analysis Evaluation
In line with the aforementioned critiques and requirements with respect to sustainability topic analysis, ESGSenticNet is tested against baselines. These include, topic models that are algebraic – NMF [30], probabilistic – LDA [9], neural – ProdLDA [52], TSCTM [60], BERTopic [21], ECRTM [59], and other publicly available sustainability dictionaries – Baier [6] Naiara [44]. We aim to answer the following research questions to assess whether the tested methods provide meaningful topic analysis for sustainability stakeholders:
  • RQ1: To what extent are the topic terms yielded by a method related to ESG, and by extension, sustainability analysis?
  • RQ2: To what extent are the ESG topic terms yielded by a method diverse, allowing analysts to view a variety of sustainability topics?
  • RQ3: To what extent do the topic terms yielded by a method express an action toward ESG, or a sustainability effort, thereby providing more material insights?
To address these research questions, topic analysis methods are evaluated according to the following.
  • ESG relatedness: The proportion of all topic terms that are ESG-related, highlighting the extent to which yielded topics are relevant to sustainability and therefore sustainability analysis.
  • Unique ESG terms: The number of unique ESG topic terms, indicating the comprehensiveness and diversity of ESG content captured.
  • ESG action orientation: The proportion of all topic terms that express an action taken toward ESG, or in other words, a sustainability effort. This indicates whether the terms are able to express information material, and of interest to stakeholders.
For the above metrics, GPT-4 is leveraged to classify all topic terms yielded from the models. GPT-4 distinguishes the following – 1) whether each topic term is ESG-related, 2) whether each topic term constitutes an action toward improving ESG. Thereafter, we aggregate the GPT-4 classification results to produce scores for ESG relation, unique ESG terms, ESG action. We validate using GPT-4 in our evaluation through an GPT4-human agreement study (section 6.6). Full experimental details for topic analysis evaluation are described in section 6.3, with the prompts provided in figure 7 (ESG relatedness), and figure 8 (ESG action orientation).
[Task Description] As a sustainability expert, your task is to classify terms based on their relevance to Environmental, Social, and Governance (ESG) factors. [Response Instructions] Evaluate each term based on its content to determine its most appropriate category. If a term is explicitly related to ESG, assign the label ‘ESG’. If a term is not related to ESG or the relationship is vague, assign the label ‘non-ESG’. Ensure all responses strictly adhere to the definitions and format provided. [Response Format] Fill in your response in the triple backticks below using the following format: Input: <term> Response: <ESG or non-ESG>
Figure 7: Prompt for evaluating ESG relatedness
[Task Description] As a sustainability expert, your task is to classify terms based on whether they express an action taken toward ESG. [Response Instructions] Evaluate each term based on its content to determine whether it describes an action taken to advance Environmental, Social, and Governance (ESG) goals. If a term describes an action taken to enhance ESG objectives, assign the label ‘True’. If a term does not describe an action taken toward ESG objectives, assign the label ‘False’. Ensure all responses strictly adhere to the definitions and format provided. [Response Format] Fill in your response in the triple backticks below using the following format: Input: <term> Response: <True or False>
Figure 8: Prompt for evaluating ESG action orientation
6.3 Experimental Settings
Topic analysis is performed on 319 sustainability disclosures from 75 large Singaporean Companies from 2015 to 2023, obtained from the respective company websites. These reports were unused during the development of ESGSenticNet. Algebraic, Probabilistic & Neural methods such as NMF [30], LDA [9], ProdLDA [52], BERTopic [21], TSCTM [60], ECRTM [59], are unsupervised, and we run them across a wide range of hyperparameters, to obtain the best results for each method. These parameters include every combination for vocab size=$=\{2500,5000,7500,10000,12500\}$, number of topics=$=\{10,20,30,40,50,60,70\}$, number of topic words=$=\{5,10,15,20,25,30\}$. Additionally, these models were deployed through the TopMost package [61]. Dictionary methods comprise terms that are categorised according to specific sustainability topics. A topic’s occurrence is computed from the frequency of occurrence of its associated terms. In our experiment, the topic terms of the dictionaries correspond to dictionary terms that appear within the corpora of sustainability disclosures. To the best of our knowledge, dictionary methods in the published literature are limited, with only a few works that make their full dictionaries publicly available. These include Baier [6] and Naiara [44].
ESGSenticNet comprises knowledge triplets in the format (concept, relation, category), delineating how specific concepts related with sustainability categories or topics. To enhance semantic specificity and context, we employ three word concepts, exclusively using those that possess the ‘supports’ relation toward a sustainability topic. Similar to dictionary methods, a topic’s occurrence is determined by the frequency of appearance of concepts that support the topic (we explore this further in a case study in section (6.5)). The topic terms of ESGSenticNet correspond to these concepts that appear within the sustainability disclosures corpora. Additionally, we experiment with two configurations for ESGSenticNet: Exact involves exact string matches of each concept, Flexible involves taking advantage of the (verb, noun phrase) arrangement of concepts for matching. To elaborate on this, the verb (1st word) is matched separately from the corresponding noun-phrase (2nd & 3rd words) within each sentence. Given the example “We reduce our water consumption”, ‘reduce’ is matched as the verb, and ‘water consumption’ is matched as the noun-phrase, with ‘reduce water consumption’ returned as the matched concept without requiring its exact string match. Flexible enables greater detection of ESG concepts by allowing for semantic variability (see study in section 6.7), while Exact is useful for detecting precise ESG impacts.
6.4 Results & Discussion
Baseline Best ESG-unique Best ESG-rel Best ESG-act
ESG-unique ESG-rel ESG-act ESG-unique ESG-rel ESG-act ESG-unique ESG-rel ESG-act
NMF 332* 0.28* 0.08* 3 0.42 0.00 332* 0.28* 0.08*
LDA 23* 0.31* 0.01* 3 0.36 0.00 23* 0.31* 0.01*
ProdLDA 334* 0.26* 0.07* 17 0.38 0.10 13 0.38 0.16
BERTopic 76 0.33 0.05 13 0.40 0.04 13 0.35 0.05
TSCTM 359 0.15 0.06 11 0.32 0.06 23 0.31 0.15
ECRTM 440 0.18 0.09 23 0.46 0.08 182 0.31 0.15
Baier 212* 0.49* 0.29* 212* 0.49* 0.29* 212* 0.49* 0.29*
Naiara 1* 0.50* 0.50* 1* 0.50* 0.50* 1* 0.50* 0.50*
ESGSenticNet (exact) 359* 0.76* 0.84* 359* 0.76* 0.84* 359* 0.76* 0.84*
ESGSenticNet (flexible) 2555* 0.79* 0.81* 2555* 0.79* 0.81* 2555* 0.79* 0.81*
Table 10: Evaluation of Topic Analysis Methods, through the metrics of unique ESG terms (ESG-unique), ESG relatedness (ESG-rel), and ESG action orientation (ESG-act). Best results are marked in bold and the best baseline results are underlined. (*) indicates the results from the same run (i.e. same parameters). Best ESG-unique, Best ESG-rel, and Best ESG-act represent the top-performing results from the respective runs of the tested methods.
Our experimental results (table 10) indicate that ESGSenticNet scores higher than the other topic analysis methods on ESG relatedness and ESG action orientation by at least 26% and 31% respectively. Additionally, both ESGSenticNet (exact) and ESGSenticNet (flexible) extract a high number of unique ESG terms, with ESGSenticNet (flexible) scoring the highest amongst all methods tested. Although requiring limited computational resources as a lexical method, ESGSenticNet produces significantly higher performance compared to the more expensive unsupervised topic modelling methods. This is despite how the topic models were exhaustively ran on numerous permutations of parameters (section 6.3) that required extensive computational costs.
Unsupervised topic models such as TSCTM, ECRTM, ProdLDA, NMF can extract a high number of unique ESG terms ($>300$). Yet, for the same runs, the concentration or prevalence of ESG terms within each topic is modest, as highlighted by their limited scores in ESG relatedness ($<0.30$). Therefore, despite capturing diverse ESG topic terms, these methods fall short in generating topics that are richly centered on ESG themes. From table 11, we highlight examples of the topics generated by these methods in comparison to ESGSenticNet, highlighting how most of the topics generated by these topic models are not ESG-centric. This limitation is critical in sustainability topic analysis, given that the primary goal is to derive insights from topics that are not only inclusive of ESG terms but are also substantially focused on ESG issues.
Moreover, topic model runs that yield the highest ESG relatedness are only able to yield few unique ESG-terms ($<25$). Furthermore, for these topic model configurations, as well as for other dictionaries such as Baier and Naiara, ESG relatedness is still $\leq 0.50$. This suggests that yielded topics are still not predominantly centred around ESG themes.
Finally, most of the tested baseline methods score modestly for ESG action orientation, Naiara noticeably higher. Yet, Naiara is only able to derive $1$ unique ESG-term, which limits the significance of its analysis. These results suggest that the baseline methods may not fully capture information that is material to stakeholders, particularly in terms of conveying sustainability actions and efforts.
Method Topic Example Topic Words
TSCTM Topic 1 offshore, baker, vessels, oil, cho, vessel, chartering, ship, sea charter
Topic 2 rsp, venture, johor, black, sites
Topic 3 controlling, provision, shall, period, accounting, mainly, listing, media, entity, time
ECRTM Topic 1 minerals, healthcare, disa, teh, bhd holdings, agm, sdn, heatec, sri
Topic 2 banks, grand employees, duty, palm
Topic 3 disa, shares, theft share, anniversary, lau, purchase, options, june cancellation
ProdLDA Topic 1 financial, value, non assets, december
Topic 2 loss, company, board, statements, directors
Topic 3 greenhouse, mountain, aspires, gresb, honeyness, feel, strategy, board, candid, phenomenon, conforms, performance, cubic, cabling, unoccupied
NMF Topic 1 duly, assets, transparency, value, priority
Topic 2 property, fair, rate, million, new
Topic 3 esg, administration, confirmed, quarter, save, present, businesses, employees, workplace, flow, focusing, ceo, default, meetings, termination
ESGSenticNet Topic 1 abide local law, achieve regulatory compliance, fulfill regulatory requirement, follow international law, comply environmental rule
Topic 2 increase employee satisfaction, improve staff engagement, invest career development, maintain employee safety
Topic 3 minimise carbon footprint, monitor ghg emission, reduce environmental pollution, lower air emission
Table 11: Examples of topic words for topic models with the highest numbers of unique ESG terms, in comparison with ESGSenticNet for qualitative analysis
6.5 Case Study - Insights offered by ESGSenticNet
Figure 9: Frequency of ESGSenticNet topics from the flexible run
Figure 10: Top ESGSenticNet concepts from the flexible run, in support of the sub topics (dark blue), and their corresponding broad topic (light blue)
The topic analysis conducted by ESGSenticNet in section 6 is visualised with the figures 9 & 10, which represent different ways to analyse the text corpus. Figure 9 provides a higher level breakdown of the topics present in the corpus. It highlights the top 5 frequently occurring broad sustainability topics present in the corpus, as well as the proportion of their constituent sub topics, in accordance with ESGSenticNet’s hierarchical taxonomy. The topic frequencies are denoted by the occurrence of matched concepts that ‘support’ a topic, leveraging ESGSenticNet’s predefined relations. Separately, figure 10 provides a more granular analysis of the corpus, highlighting matched concepts that explicitly convey sustainability actions, thereby communicating important sustainability information. Figure 10 displays the top occurring concepts that support the topics shown in figure 9. Put together, figures 9 and 10 highlight different methods to analyse a sustainability corpus, which can be utilised depending on a stakeholder’s interest and preference.
6.6 GPT-4 Human Expert Agreement on Topic Analysis Evaluation Metrics
GPT-4 evaluations have shown significant reliability in the literature [64]. A study is conducted to assess GPT-4’s ability for the following tasks– i) classifying topic terms as ESG-related, ii) classifying topic terms as actions taken toward ESG. A random sample of 510 topic terms were gathered for each task. GPT-4 annotations for each set of topic terms are compared with annotations from 2 different groups of 3 human experts each, with each group handling a different task. An average agreement score of 83.7% observed for task i) and 91.8% for task ii).
For evaluating GPT-4 human agreement, the following instructions were given for classifying topic terms as ESG-related.
The table below presents various terms. Determine if each term is ESG-related. Mark ‘TRUE’ if the term is ESG-related and ‘FALSE’ otherwise.
For evaluating GPT-4 human agreement, the following instructions were given for classifying topic terms as an action taken toward ESG.
The table below presents various terms. Determine if each term expresses an action toward improving a company’s ESG performance. Mark ‘TRUE’ if the term express an action toward improving ESG performance and ‘FALSE’ otherwise.
6.7 Effectiveness of ESGSenticNet Flexible Matching Method
A study is conducted to assess the precision of matching ESGSenticNet concepts through the ‘flexible’ method. A random sample of 200 sentences matched through this method was collected, excluding sentences matched via the ‘precise’ method. Human annotators were tasked with evaluating whether each sentence clearly conveys the idea expressed represented by its matched concept. From our results, 75.0% of sentences comprised the idea expressed by its corresponding matched concepts. This suggests that the ‘flexible’ method retains a moderately high level of precision, despite allowing for greater semantic variability. For evaluating the ESGSenticNet flexible matching method, the following instructions were given to evaluate if each sentence clearly conveys the idea expressed by its matched concept. The table below presents different sentences and a corresponding concept. Determine if each sentence clearly conveys the idea expressed by its matched concept. Mark ‘TRUE’ if the sentence clearly conveys the concept’s idea, and ‘FALSE’ otherwise.
7 Conclusion & Future Work
This paper suggests that by focusing on the identified challenges of immateriality, complexity and subjectivity, NLP methods can be enhanced for sustainability analysis. As our findings indicate, ESGSenticNet, developed with these challenges in mind, provides more relevant and actionable insights for sustainability stakeholders, presentiing a significant improvement over existing sustainability lexicons, let alone state-of-the-art NLP methods for corporate sustainability analysis. This is despite how ESGSenticNet’s deployment does not entail computational training or technical expertise. Ultimately, ESGSenticNet aims to be a tool that can be readily used by stakeholders regardless of their technical background, to democratise access to crucial corporate sustainability insights.
Appendix A Additional details on Human Validation
For evaluating ESGSenticNet accuracy (section 5.9), ESGSenticNet’s flexible matching method (section 6.7), GPT-4 human agreement (section 6.6), there are a total of 9 volunteer human annotators. These annotators are based in Singapore, comprising researchers from the Asian Institute of Digital Finance, Sustainable and Green Finance Institute, who are pursuing doctoral-level and post-doctoral research within the corporate sustainability field. All human validation studies follow the same annotation scheme. When there are disagreements, discussions are held to achieve a resolution. Where there is no clear resolution, majority voting is taken to determine the ground truth label. For the derivation of the taxonomy for ESGSenticNet, we enlisted the help of two tenured researchers from the Singapore Sustainable and Green Finance Institute (SGFIN), whom we deem as experts within the sustainable finance field.

### v2

@@
 Abstract
-Evaluating corporate sustainability performance is essential to drive sustainable business practices, amid the need for a more sustainable economy.
-However, this is hindered by the complexity and volume of corporate sustainability data (i.e.
-sustainability disclosures), not least by the effectiveness of the NLP tools used to analyse them.
-To this end, we identify three primary challenges – immateriality, complexity, and subjectivity, that exacerbate the difficulty of extracting insights from sustainability disclosures.
-To address these issues, we introduce ESGSenticNet, a publicly available knowledge base for sustainability analysis.
-ESGSenticNet is constructed from a neurosymbolic framework that integrates specialised concept parsing, GPT-4o inference, and semi-supervised label propagation, together with a hierarchical taxonomy.
-This approach culminates in a structured knowledge base of 44k knowledge triplets – (‘halve carbon emission’, supports, ‘emissions control’), for effective sustainability analysis.
-Experiments indicate that ESGSenticNet, when deployed as a lexical method, more effectively captures relevant and actionable sustainability information from sustainability disclosures compared to state of the art baselines.
-Besides capturing a high number of unique ESG topic terms, ESGSenticNet outperforms baselines on the ESG relatedness and ESG action orientation of these terms by 26% and 31% respectively.
-These metrics describe the extent to which topic terms are related to ESG, and depict an action toward ESG.
-Moreover, when deployed as a lexical method, ESGSenticNet does not require any training, possessing a key advantage in its simplicity for non-technical stakeholders.
+Identifying Environmental, Social, and Governance (ESG) topics in sustainability reports, or ESG topic analysis, is essential for assessing a company’s ESG performance.
+While commonly used for this task, existing NLP approaches face three key challenges: irrelevance, the extracted topics are unrelated to ESG, immateriality, the extracted topics lack meaningful value, and limited organisation, the extracted topics have poor alignment with ESG frameworks.
+To address these challenges, we introduce ESGSenticNet, a publicly available knowledge base built from 1,679 SGX sustainability reports.
+ESGSenticNet is constructed using a neurosymbolic framework—combining symbolic reasoning (structured parsing rules) with neural language models (GPT-4o)—to identify and categorise ESG-related concepts.
+The process begins by extracting concepts (or meaningful phrases) from reports using parsing methods, which are then labelled according to a hierarchical ESG taxonomy using GPT-4o.
+These labelled concepts are further expanded using a graph-based method that propagates labels to similar concepts without requiring additional manual annotation.
+The final knowledge base consists of 44,232 structured entries, or “triples,” in the format (concept, relation, ESG topic)—e.g., (“halve carbon emission”, supports, “emissions control”)—which serve as ESG lexicons for topic analysis.
+This allows ESG topics to be identified by matching concepts found in reports to their associated topics.
+We evaluate ESGSenticNet on 319 sustainability reports against existing NLP methods.
+ESGSenticNet outperforms baselines, producing topic terms with 76% ESG-relatedness (+26% improvement) and 84% ESG action-orientation (+34% improvement), while capturing a high number of unique ESG terms (359).
+Human evaluation of 500 concepts shows that ESGSenticNet achieves 86%+ accuracy in categorising concepts under correct ESG topics.
+This suggests that during topic analysis, ESGSenticNet shows a strong degree of reliability in aligning topic terms with their intended topics.
+Additionally, ESGSenticNet’s deployment requires no training, making it accessible for stakeholders regardless of resources or expertise.
+Our key contributions include: (1) developing a neurosymbolic framework for deriving sustainability lexicons, which can be potentially extended to other contexts such as climate risk disclosures; (2) introducing new metrics for ESG topic analysis, which we argue are more meaningful for the task at hand; and (3) releasing ESGSenticNet as a public tool that effectively addresses the core challenges of irrelevance, immateriality, and limited organisation in ESG topic analysis.
 keywords:
 Generative AI, Large Language Models, Knowledge Base, Lexicons, Corporate Sustainability Analysis
-journal: Information Processing and Management
+journal: npj Artificial Intelligence
 […]
@@
 […]
-[8]organization=Honorary Research Fellow, Imperial College London,
+[8]organization=Honorary Research Fellow, Imperial College London, addressline=Exhibition Road, London SW7 2AZ, country=United Kingdom
 1 Introduction
-With growing concerns over social inequality and global warming, it is imperative to ensure the sustainability of public companies [13].
-To this end, accurate insights into the sustainability efforts of corporations are crucial to drive responsible business practices [40].
-However, the complexity and volume of corporate sustainability data [38], often disclosed through sustainability reports (i.e.
-sustainability disclosures), can make these insights inaccessible without specialised tools.
-While NLP tools can unlock insights from this vast array of corporate sustainability data, their deployment remains nascent within this field, often lacking the specialised design needed for corporate sustainability analysis.
-Specifically, they do not address the challenges of immateriality, complexity and subjectivity, which are frequently associated with sustainability data [41].
-  • Immateriality: Sustainability disclosures are lengthy, containing significant amounts of irrelevant content for sustainability analysis [51, 28].
-Key or material information for stakeholders can often be obfuscated [49], making their extraction via NLP tools challenging.
-  • Complexity: Sustainability ideas are sophisticated, and are typically expressed by complex syntactic constructions [50].
-The distinctiveness in sustainability language necessitates that NLP tools must encompass more than just general language understanding capabilities.
-  • Subjectivity: Stakeholders are interested in diverse sustainability areas, and utilise different frameworks and methods for sustainability analysis [39, 46].
-Therefore, designing an NLP tool that comprehensively meets the needs of different stakeholders is challenging.
-As a result, NLP algorithms, including the state of the art, are largely ineffective at extracting useful corporate sustainability insights.
-In our experiments later in this paper (section 6.3), we validate this ineffectiveness by benchmarking these algorithms in the sustainability topic analysis task.
-In response to these limitations, we develop ESGSenticNet, a publicly available knowledge base that comprises 44k sustainability knowledge triplets from 23k unique concepts, specifically designed to mine insights from sustainability disclosures.
-To address the aforementioned issues of complexity, immateriality and subjectivity, ESGSenticNet is constructed from a neurosymbolic framework that blends symbolic linguistics rules with sub-symbolic reasoning through deep learning.
-The symbolic component of this framework involves specialised concept parsing that leverages predefined linguistic patterns and rules within the sustainability domain to identify meaningful concepts within sustainability text, addressing the complexity of sustainability language.
-Additionally, these rules focus on the most pertinent linguistic structures in sustainability, enabling the identification of key sustainability information such as sustainability efforts and actions to reduce immateriality [7].
-Building on the structured outputs of the symbolic component, the subsymbolic component integrates GPT-4o inference with semi-supervised label propagation to classify the relation between the parsed concepts and their respective sustainability categories.
-By interpreting nuanced language variations and contextual subtleties that symbolic rules cannot fully capture, this component ensures that concepts are organised according to key sustainability categories, further enhancing their materiality.
-Additionally, these categories within ESGSenticNet are arranged according to a hierarchical and comprehensive taxonomy, addressing subjectivity by allowing concepts to be leveraged at different levels of granularity, tailored to the specific interest of stakeholders.
-The result is an extensive knowledge base, ESGSenticNet, that comprises actionable information in the form of knowledge triplets.
-These triplets highlight concepts and their relation toward sustainability themes, such as (‘minimise paper waste’ (concept), supports (relation), ‘waste management’ (category)).
-We evaluate ESGSenticNet’s ability to extract meaningful insights from sustainability disclosures, by deploying the knowledge base as a structured lexicon method for sustainability topic analysis.
-In this task, the concepts within ESGSenticNet’s knowledge triplets are matched from a large sustainability corpora.
-ESGSenticNet improves on established baselines significantly, more extensively capturing sustainability related information, including material content that highlight sustainability efforts.
-ESGSenticNet yields a significant number of ESG topic terms, outperforming existing state of the art baselines by at least 26% and 31% respectively on the ESG relatedness and ESG action orientation of these terms.
-This highlights demonstrates ESGSenticNet’s potential to contribute meaningfully to sustainability analysis.
-Moreover, as a lexical method, ESGSenticNet is able to extract ESG topic terms without any training or significant technical expertise.
-This significantly allows it to be leveraged by stakeholders of different backgrounds.
+Amid growing demands for corporate sustainability—companies’ commitment to reducing environmental impact, promoting social responsibility, and upholding ethical governance [54]—regulators and investors are increasingly focused on corporate sustainability analysis [16].
+This involves evaluating a company’s progress toward Environmental, Social, Governance (ESG) goals, otherwise known as ESG assessment, with company-disclosed sustainability reports often used for analysis [52].
+However, the complexity and sheer volume [48] of these reports typically render manual analysis impractical [53].
+To overcome this challenge, NLP-driven ESG topic analysis — the automatic identification of ESG topics within sustainability reports — has been widely implemented.
+By enabling the automated and structured analysis of sustainability reports, ESG topic analysis provides regulators and investors key insights on a corporation’s sustainability performance, such as extent of alignment with environmental objectives [70], future orientation of sustainability activities [34], among others.
+These insights are useful for a variety of applications, such as sustainable portfolio construction [35] and regulatory auditing [27], which significantly underpin the field of sustainable finance.
+However, despite advancements in NLP, existing methods struggle to meet the specialised demands of ESG topic analysis, limiting their effectiveness for this task.
+ESG topic analysis often aims to capture ESG-related content that is both relevant for ESG assessments and aligned with established ESG frameworks [54], such as SASB [25] and GRI [43].
+To accomplish this, commonly used NLP techniques, including rule-based dictionary methods [5] and unsupervised topic modeling [10], may fall short as they lack task-specific design.
+We identify three key shortcomings below:
+  • Irrelevance: Existing topic modeling methods, while excellent at producing topic terms without supervision [1], cannot be tuned to capture ESG content.
+Therefore, captured topic terms may not be relevant to ESG, leading to topics irrelevant for ESG analysis.
+Additionally, sustainability dictionaries that are popularly adapted for ESG topic analysis - i.e Baier [37], comprise a considerable portion of non-ESG terms.
+  • Immateriality: Stakeholders often prioritise content that explicitly conveys the ESG efforts or actions undertaken by corporations, as this content has greater materiality or significance for ESG performance [8].
+However, both rule-based dictionaries and topic models do not focus on extracting topic terms that capture ESG-oriented actions.
+To better align with stakeholder interests, topic terms should be refined to reflect these actions.
+  • Limited Organisation: Due to its unsupervised nature, topic modeling may generate topics that are not organised according to ESG frameworks.
+Consequently, even if ESG-related topics emerge, their lack of organisation means that they fail to comprehensively capture all dimensions of ESG.
+This incomplete coverage can be problematic for sustainability stakeholders, who rely on structured ESG frameworks to ensure that all relevant ESG aspects are analysed [44, 26].
+While sustainability dictionaries offer a more organised approach [34], their limited size (fewer than 800 terms) constrains their coverage, potentially restricting the range of ESG topics identified in practice.
+Addressing the above limitations requires an ESG topic analysis approach that explicitly captures relevant, material, and well-structured ESG content.
+To this end, we explore the following research questions:
+  • RQ1: How can a topic analysis method be designed to mitigate the limitations of existing NLP techniques in capturing relevant and material ESG content according to structured ESG frameworks?
+  • RQ2: What criteria and evaluation methods are appropriate for assessing the effectiveness of ESG topic analysis?
+  • RQ3: To what extent does a method specifically designed for ESG topic analysis outperform existing NLP approaches for this task?
+To contribute to these research questions, we develop ESGSenticNet, an NLP tool specifically designed to facilitate ESG topic analysis.
+We then evaluate it extensively against state-of-the-art NLP methods, using specialised evaluation metrics tailored for ESG topic analysis.
+ESGSenticNet is a publicly available knowledge base comprising 44k sustainability knowledge triples from 23k unique concepts.
+While ESGSenticNet itself is a knowledge base, it functions as a structured lexicon for rule-based topic analysis.
+To elaborate, each triple within ESGSenticNet follows the format (concept, relation, ESG topic)— i.e.
+(‘minimise paper waste’, supports, ‘waste management’).
+During topic analysis, the concepts serve as topic terms, which can be matched within sustainability reports to indicate the presence of specific ESG topics.
+Unlike existing topic analysis methods, ESGSenticNet explicitly addresses irrelevance, immateriality and limited organisation, through its construction from a neurosymbolic framework.
+This framework integrates symbolic linguistic rules and sub-symbolic deep-learning to ensure ESGSenticNet’s triples comprise relevant and material ESG concepts that are effectively organised according to structured ESG topics.
+To reduce irrelevance, the symbolic component applies sustainability-specific linguistic patterns to extract ESG-related concepts, thereby minimising the occurrence of irrelevant topic terms in ESG topic analysis.
+To address immateriality, the applied linguistic patterns also prioritise extracting concepts that explicitly describe ESG actions, making topic terms more meaningful for stakeholders seeking actionable ESG insights.
+To ensure organisation, the sub-symbolic component leverages GPT-4o inference, supported with a graph-based semi-supervised method, to categorise concepts according to structured ESG topics under our constructed taxonomy.
+Put together, the neurosymbolic framework yields triples in the form of (concept, relation, ESG topic), and ensures that ESGSenticNet addresses the shortcomings of existing NLP for ESG topic analysis.
+We evaluate ESGSenticNet’s effectiveness in ESG topic analysis by benchmarking it against state-of-the-art NLP algorithms commonly used for this task.
+Given the unique requirements of ESG topic analysis—capturing ESG content material for assessments, and aligned with established ESG frameworks—we introduce specialised evaluation metrics, namely (1) ESG-relatedness: The proportion of all topic terms relevant to ESG; (2) ESG-action orientation: The proportion of all topic terms focused on ESG actions, representing their material significance in sustainability assessments; (3) Unique ESG topic terms: The number of unique ESG terms, which highlights the comprehensiveness and diversity of ESG content captured.
+(1) and (2) directly addresses the core challenges of irrelevance and immateriality respectively, and (3) is related to the method’s practical utility in real-world deployments.
+We show that ESGSenticNet significantly outperforms existing baselines on these three metrics, achieving 76% ESG-relatedness (+26% improvement) and 84% ESG action-orientation (+34% improvement), while capturing a high number of unique ESG terms 359 under the exact implementation; 2555 under the flexible implementation).
+Additionally, human evaluation confirms that ESGSenticNet correctly categorises concepts under their respective ESG topics with 86%+ accuracy.
+As a result, during topic analysis, the topic terms matched from ESGSenticNet align with their associated ESG topics with considerable reliability.
+Finally, as a lexical method, ESGSenticNet requires no training or technical expertise, making it accessible to stakeholders across diverse backgrounds.
+To underscore the significance of these results for ESG topic analysis, we relate them back to the three core challenges—irrelevance, immateriality, and limited organisation—as well as the practical concerns of comprehensiveness & diversity and computational complexity.
+While the three core challenges have been previously defined, comprehensiveness & diversity captures the extent and variety of ESG content each method is able to extract, and computational complexity refers to the ease of implementation and resource demands of each method.
+Table 1 qualitatively compares ESGSenticNet, topic models, and existing lexicons in terms of how well they address each of these five dimensions.
+These assessments are informed by our experimental results—ESG-relatedness corresponds to irrelevance, ESG action-orientation to immateriality, and the number of unique ESG terms to comprehensiveness & diversity—as well as qualitative evaluations for limited organisation and computational complexity.
+[…]
+Topic Models Extracted terms are largely irrelevant to ESG topics Extracted terms rarely reflect concrete ESG action-orientation, reflecting immateriality Unstructured and not aligned with ESG frameworks Specific models capture a high amount of ESG terms, although majority of extracted terms are not ESG related High: requires extensive hyperparameter tuning and model selection
+Existing Lexicons More relevant than topic models, but still often include unrelated terms Partially captures but still largely omits ESG action-oriented terms Structured by a single ESG taxonomy Captures a small to moderate amount of ESG terms Low: no training or tuning required
+[…]
+Table 1: Qualitative comparison of ESG text analysis methods across five key dimensions.
+Lim-Org stands for limited organisation, Compre-Div for comprehensiveness & diversity, Complexity for computational complexity.
+Each assessment reflects evidence from our experimental findings where applicable — for example, high irrelevance corresponds to low ESG-relatedness scores, high immateriality to low ESG action-orientation scores, and limited comprehensiveness to low unique ESG terms captured.
+Limited organisation and computational complexity are assessed through qualitative evaluations.
+The findings position ESGSenticNet as a tool that significantly enhances ESG topic analysis, offering clear advantages over existing approaches across all five dimensions.
 The rest of our paper is arranged as follows.
-We first develop a comprehensive and hierarchical sustainability taxonomy for organising our concepts (section 4).
-Secondly, we delve into the extraction of concepts via ESG Concept Parser (section 5.2), processing the concepts to ensure coherence (section 5.3), as well as our methodology for classifying these concepts via GPT-4o and semi-supervised label propagation (sections 5.45.55.65.7).
-Finally, we evaluate the accuracy of ESGSenticNet labels (section 5.9), before running extensive experiments to assess the effectiveness of ESGSenticNet for sustainability analysis against established baselines (section 6).
-The development process for ESGSenticNet is summarised by figure 1, showcasing the key steps for concept parsing and taxonomy derivation, and how they are intertwined with the labelling process (1-4).
-2 Literature Review
-NLP-driven Sustainability Analysis.
-To enhance sustainability analysis, NLP can be deployed for different tasks, such as analysing sustainability topics [65], sustainability sentiment classification [43], extracting textual patterns [8], among other tasks.
-The insights extracted from NLP can powerfully reduce time costs and increase the consistency of evaluation [2, 15], making NLP a useful tool for sustainability analysis.
-Yet, despite NLP’s potential applications, it has only been explored within sustainability analysis to a limited degree, often lacking specialised design [41].
-This motivates the development of specially designed NLP tools for corporate sustainability analysis.
-Lexicons for Sustainability Analysis.
-For corporate sustainability analysis, lexical approaches have been widely leveraged for textual analysis [22], due to their simplicity and relevance to financial text analysis [53, 55].
-To facilitate lexical analysis, sustainablity-specific lexicons have been developed.
-Approaches such as Baier’s dictionary [6], have focused on developing single-word lexicons according to ESG pillars.
-However, while Baier’s dictionary has been a popular choice for sustainability analysis works [25], single word lexicons cannot adequately capture the semantics of the sustainability field.
-In contrast, dictionaries that have collected sustainability related text spans (of multiple-grams) [17, 29], are limited, not publicly available, and not comprehensively labelled beyond ESG pillars.
-Additionally, due to their reliance on manual human-annotation, these dictionaries have an exceedingly small number of lexicons (less than 800), raising questions on their comprehensiveness, and consequently their effectiveness for capturing textual patterns.
-This motivates the development of an extensive knowledge base that can function as a structured dictionary of concepts, and the exploration of automated methods such as NLP for constructing and organising sustainability lexicons.
-Knowledge Base Construction.
-Methods for constructing knowledge bases can vary according to the domain and application [26].
-Notable knowledge bases, such as ConceptNet [31] involved concept parsing and human annotation, SenticNet 5 [11] leveraged deep learning and lexical substitution, question answering via advanced graphical neural networks (GNNs) [63].
-Newer avenues for knowledge construction involve generative-LLMs, which have shown strong performance in related tasks such as reasoning, construction, and generalisation [67].
-Unlike all the previous methods which heavily rely on semantic similarities for labelling, generative LLMs possess powerful emergent capabilities that extend beyond textual similarities [58].
-These include reasoning, knowledge relationship construction, and generalisation to different domains [67].
-Moreover, unlike GNNs that rely on heavily structed text data [32], generative-LLMs can flexibly adapt to unstructured text.
-This, coupled with added domain flexibility and inferencing capabilities on limited textual cues, motivates our use of generative-LLMs for knowledge derivation within sustainability.
-3 Data
-1680 sustainability disclosures from a wide range of Singapore Exchange (SGX) based companies, in the period of 2017 to 2022 inclusive, are utilised for taxonomy and knowledge base construction.
-These reports can be found online via respective company websites, and on request, we can provide a list of the reports utilised (i.e.
-the companies and year of reporting).
-Figure 1: Overview of ESGSenticNet Development
-4 Sustainability Taxonomy
-Firms may not report in accordance with existing sustainability taxonomies (i.e.
-GRI [33]), given that their operations are complex and may not be adequately captured by these taxonomies [41].
-To further align taxonomies with sustainability report content, data-driven approaches can be utilised for taxonomy derivation [5].
-Yet, empirical approaches in the literature are limited, and these methods do not involve human-in-the-loop inference to ensure the meaningfulness of constructed sustainability topics.
-These challenges motivate our taxonomy construction, that, while empirically driven, involves human-in-the-loop inference to ensure the relevance and meaningfulness of categories.
-4.1 Methodology
-Sustainability disclosures are tokenised into separate sentences, with each sentence transformed into embeddings by Sentence-BERT (S-BERT) [45].
-BERTopic is then run on these embeddings, with the configuration of KeyBERT [20].
-The topic words are then provided to sustainability experts (A) to facilitate their development of the taxonomy for ESGSenticNet (table 1).
-Examples of topic words and their inferred topics can be found in table 1.
-The result is a taxonomy (table 2) with the following features.
-  • Realistic: Given that taxonomy development leverages empirical insights from the sustainability reports themselves, categories can capture content within sustainability reports that might otherwise be overlooked by standard frameworks.
-For instance, ‘Green Technology’ is derived as a taxonomy category, while not being explicitly emphasised in standard reporting frameworks (i.e.
-GRI [33]).
-  • Activities Cross-Theme: To capture the interconnectedness of sustainability dimensions [38], categories include areas of a firm’s activities (‘operations’, ‘development’ etc.), on top of sustainability impacts (‘emissions control’, ‘waste management’ etc.).
-In contrast with existing frameworks, this offers insight into the sustainability of specific business activities.
-  • Granularity: The taxonomy has a hierarchical nature to enable analysis at different levels of granularity.
-Topic keywords Category
+We begin with a literature review (Section 5).
+Then, we explain the preliminary steps (Section 4) prior to constructing the knowledge base, which involves describing the data utilised in this study (Section 4.1), and defining the knowledge base’s organisational framework.
+The latter involves establishing a hierarchical ESG taxonomy that defines the topics in our knowledge base (sections 4.24.3) and specifying the relations and rules that govern its structure (Section 4.4).
+After outlining the prelimnary steps, we delve into our neurosymbolic framework for constructing the knowledge triples within ESGSenticNet (Section 4.5).
+This involves the extraction of concepts via ESG concept parser (Section 4.6), quality control—processing the extracted concepts to ensure their coherence (Section 4.7), followed by the labelling process for classifying these concepts according to our taxonomy topics (Section 4.8).
+This labelling process involves leveraging GPT-4o inference, supported by a graph-based semi-supervised method (sections 4.8.14.8.24.8.34.8.4).
+Our neurosymbolic framework (Section 4.5) results in the construction of ESGSenticNet, which we extensively evaluate through experiments on topic analysis, with the results and implications discussed (Section 2).
+2 Results
+To evaluate the effectiveness of ESGSenticNet against existing NLP methods, we perform topic analysis on the evaluation set (described in Section 4.1).
+This set comprises 319 sustainability reports from SGX companies, and the reports were unused during the development of ESGSenticNet.
+In the following, we discuss the evaluation metrics (Section 2.1) and experimental settings (Section 2.2) utilised during this experiment, before discussing the experimental results (Section 2.3).
+Thereafter, we further evaluate the accuracy of ESGSenticNet’s triples through human evaluation (Section 2.4), before discussing a case study of ESGSenticNet’s application (Section 3.1) and the implications of our ESGSenticNet study (Section 3.2).
+We finish this section with additional ablation studies (Section 4.10) to improve the robustness of our findings.
+2.1 Evaluation Metrics
+As discussed earlier (Section 1), ESG topic analysis has distinct requirements as it supports sustainability assessments.
+It focuses on extracting ESG content that is significant for these assessments, and organised within established ESG frameworks.
+Therefore, while topic analysis metrics—specifically pertaining to topic models (i.e.
+perplexity, PMI) [1]—have been explored, ESG topic analysis require specialised metrics.
+Therefore, methods are evaluated according to the following:
+  • ESG relatedness: The proportion of all topic terms that are ESG-related, highlighting the extent to which yielded topics are relevant to ESG and therefore ESG analysis.
+  • ESG action orientation: The proportion of all topic terms that express an action taken toward ESG, or in other words, an ESG effort.
+This indicates whether the terms are able to express information material to ESG assessments, and of interest to stakeholders.
+  • Unique ESG terms: The number of unique ESG topic terms, indicating the comprehensiveness and diversity of ESG content captured.
+For the above metrics, GPT-4 is leveraged to classify all topic terms yielded from the models, aligning with the LLM-as-a-judge method from established NLP studies [80].
+GPT-4 distinguishes the following – 1) whether each topic term is ESG-related, 2) whether each topic term constitutes an action toward improving ESG.
+Thereafter, we aggregate the GPT-4 classification results to produce scores for ESG relation, unique ESG terms, ESG action.
+We utilise GPT-4 in our work given that it offers a scalable, automated, and reproducible approach to evaluating ESG topic terms [18] — a task that would be prohibitively costly and impractical to perform manually, given the tens of thousands of topic terms involved.
+At the same time, we recognise the potential limitations of LLM-based evaluation [80].
+Therefore, to ensure robustness, we validate the use of GPT-4 in our evaluation through an GPT4-human agreement study (Section 4.10.2).
+Full experimental details for topic analysis evaluation are described in Section 2.2, with the prompts provided in Figure 1 (ESG relatedness), and Figure 2 (ESG action orientation).
+[Task Description] As a sustainability expert, your task is to classify terms based on their relevance to Environmental, Social, and Governance (ESG) factors.
+[Response Instructions] Evaluate each term based on its content to determine its most appropriate topic.
+If a term is explicitly related to ESG, assign the label ‘ESG’.
+If a term is not related to ESG or the relationship is vague, assign the label ‘non-ESG’.
+Ensure all responses strictly adhere to the definitions and format provided.
+[Response Format] Fill in your response in the triple backticks below using the following format: Input: <term> Response: <ESG or non-ESG>
+Figure 1: Prompt for evaluating ESG relatedness
+[Task Description] As a sustainability expert, your task is to classify terms based on whether they express an action taken toward ESG.
+[Response Instructions] Evaluate each term based on its content to determine whether it describes an action taken to advance Environmental, Social, and Governance (ESG) goals.
+If a term describes an action taken to enhance ESG objectives, assign the label ‘True’.
+If a term does not describe an action taken toward ESG objectives, assign the label ‘False’.
+Ensure all responses strictly adhere to the definitions and format provided.
+[Response Format] Fill in your response in the triple backticks below using the following format: Input: <term> Response: <True or False>
+Figure 2: Prompt for evaluating ESG action orientation
+2.2 Experimental Settings
+For this study, the following baselines were utilised.
+Topic models that are algebraic – NMF [40], probabilistic – LDA [10], neural – ProdLDA [67], TSCTM [74], BERTopic [30], ECRTM [73], and other publicly available ESG dictionaries – Baier [5] Naiara [60].
+To the best of our knowledge, ESG dictionary methods in the published literature are limited, with Baier and Naiara among the only few works that make their full dictionaries publicly available.
+[…]
+Dictionary methods comprise terms that are categorised according to ESG topics, with a topic’s occurrence computed from the frequency of occurrence of its associated terms.
+In our experiment, the topic terms of the dictionaries correspond to dictionary terms that appear within the corpora of sustainability reports.
+ESGSenticNet comprises knowledge triples in the format (concept, relation, ESG topic), delineating how specific concepts related with ESG topics or topics.
+To enhance semantic specificity and context, we employ three word concepts, exclusively using those that possess the ‘supports’ relation toward a ESG topic.
+Similar to dictionary methods, a topic’s occurrence is determined by the frequency of appearance of concepts that support the topic (we explore this further in a case study in section (3.1)).
+The topic terms of ESGSenticNet correspond to these concepts that appear within the sustainability reports corpora.
+Additionally, we experiment with two configurations for ESGSenticNet: Exact involves exact string matches of each concept, Flexible involves taking advantage of the (verb, noun phrase) arrangement of concepts for matching.
+To elaborate on this, the verb (1st word) is matched separately from the corresponding noun-phrase (2nd & 3rd words) within each sentence.
+Given the example “We reduce our water consumption”, ‘reduce’ is matched as the verb, and ‘water consumption’ is matched as the noun-phrase, with ‘reduce water consumption’ returned as the matched concept without requiring its exact string match.
+Flexible enables greater detection of ESG concepts by allowing for semantic variability (see study in Section 4.10.3), while Exact is useful for detecting precise ESG impacts.
+2.3 Experimental Results
+Baseline Best ESG-unique Best ESG-rel Best ESG-act
+ESG-unique ESG-rel ESG-act ESG-unique ESG-rel ESG-act ESG-unique ESG-rel ESG-act
+NMF 332* 0.28* 0.08* 3 0.42 0.00 332* 0.28* 0.08*
+LDA 23* 0.31* 0.01* 3 0.36 0.00 23* 0.31* 0.01*
+ProdLDA 334* 0.26* 0.07* 17 0.38 0.10 13 0.38 0.16
+BERTopic 76 0.33 0.05 13 0.40 0.04 13 0.35 0.05
+TSCTM 359 0.15 0.06 11 0.32 0.06 23 0.31 0.15
+ECRTM 440 0.18 0.09 23 0.46 0.08 182 0.31 0.15
+Baier 212* 0.49* 0.29* 212* 0.49* 0.29* 212* 0.49* 0.29*
+Naiara 1* 0.50* 0.50* 1* 0.50* 0.50* 1* 0.50* 0.50*
+ESGSenticNet (exact) 359* 0.76* 0.84* 359* 0.76* 0.84* 359* 0.76* 0.84*
+ESGSenticNet (flexible) 2555* 0.79* 0.81* 2555* 0.79* 0.81* 2555* 0.79* 0.81*
+Table 2: Evaluation of Topic Analysis Methods, through the metrics of unique ESG terms (ESG-unique), ESG relatedness (ESG-rel), and ESG action orientation (ESG-act).
+Best results are marked in bold and the best baseline results are underlined.
+(*) indicates the results from the same run (i.e.
+same parameters).
+Best ESG-unique, Best ESG-rel, and Best ESG-act represent the top-performing results from the respective runs of the tested methods.
+Figure 3: Heatmap highlighting performance of different topic analysis methods on the metrics of number of unique ESG terms (ESG-unique), ESG relatedness (ESG-rel), ESG action orientation (ESG-act).
+ESGSenticNet(Exact) refers to ESGSenticNet deployed by Exact matching, while ESGSenticNet(Flexible) refers to ESGSenticNet deployed by Flexible matching, as explained in Section 5.2.
+Darker green indicates better performance, lighter green indicates poorer performance.
+Our experimental results (Table 2) indicate that ESGSenticNet scores higher than the other topic analysis methods on ESG relatedness and ESG action orientation by at least 26% and 31% respectively.
+Additionally, both ESGSenticNet (exact) and ESGSenticNet (flexible) extract a high number of unique ESG terms, with ESGSenticNet (flexible) scoring the highest amongst all methods tested.
+Although requiring limited computational resources as a lexical method, ESGSenticNet produces significantly higher performance compared to the more expensive unsupervised topic modeling methods.
+This is despite how the topic models were exhaustively ran on numerous permutations of parameters (Section 2.2) that required extensive computational costs.
+Method Topic Example Topic Words
+TSCTM Topic 1 offshore, baker, vessels, oil, cho, vessel, chartering, ship, sea charter
+Topic 2 rsp, venture, johor, black, sites
+Topic 3 controlling, provision, shall, period, accounting, mainly, listing, media, entity, time
+ECRTM Topic 1 minerals, healthcare, disa, teh, bhd holdings, agm, sdn, heatec, sri
+Topic 2 banks, grand employees, duty, palm
+Topic 3 disa, shares, theft share, anniversary, lau, purchase, options, june cancellation
+ProdLDA Topic 1 financial, value, non assets, december
+Topic 2 loss, company, board, statements, directors
+Topic 3 greenhouse, mountain, aspires, gresb, honeyness, feel, strategy, board, candid, phenomenon, conforms, performance, cubic, cabling, unoccupied
+NMF Topic 1 duly, assets, transparency, value, priority
+Topic 2 property, fair, rate, million, new
+Topic 3 esg, administration, confirmed, quarter, save, present, businesses, employees, workplace, flow, focusing, ceo, default, meetings, termination
+ESGSenticNet Topic 1 abide local law, achieve regulatory compliance, fulfill regulatory requirement, follow international law, comply environmental rule
+Topic 2 increase employee satisfaction, improve staff engagement, invest career development, maintain employee safety
+Topic 3 minimise carbon footprint, monitor ghg emission, reduce environmental pollution, lower air emission
+Table 3: Examples of topic words for topic models with the highest scores on the number of unique ESG terms.
+Unsupervised topic models such as TSCTM, ECRTM, ProdLDA, NMF can extract a high number of unique ESG terms (>300>300).
+Yet, for the same runs, the concentration or prevalence of ESG terms within each topic is modest, as highlighted by their limited scores in ESG relatedness (<0.30<0.30).
+Therefore, despite capturing diverse ESG topic terms, these methods fall short in generating topics that are richly centered on ESG themes.
+From Table 3, we highlight examples of the topics generated by these methods in comparison to ESGSenticNet, highlighting how most of the topics generated by these topic models are not ESG-centric.
+This limitation is critical in ESG topic analysis, given that the primary goal is to derive insights from topics that are not only inclusive of ESG terms but are also substantially focused on ESG issues.
+We also compare the number of unique ESG-related terms extracted by methods that construct structured ESG lexicons from existing sustainability corpora—specifically Baier and Naiara, as this serves as an indicator of lexical coverage and diversity.
+ESGSenticNet yields 359 unique terms in its exact matching configuration and 2,555 under its more flexible variant, compared to 212 for Baier and just 1 for Naiara.
+These differences suggest that ESGSenticNet captures a broader and more diverse set of ESG terms from the test corpora, reflecting a more extensive and diverse collection of ESG lexicons (or ESG concepts) relative to the Baier and Naiara lexicons.
+This outcome reflects the advantages of our neuro-symbolic framework—discussed in Section 5.2—over manual or anchor-based lexicon construction methods, as it enables scalable and systematic identification of ESG concepts through automated concept parsing, GPT-4o-based annotation, diversity-aware seed selection, and label propagation.
+Moreover, topic model runs that yield the highest ESG relatedness are only able to yield few unique ESG-terms (<25<25).
+Furthermore, for these topic model configurations, as well as for other dictionaries such as Baier and Naiara, ESG relatedness is still $\leq 0.50$.
+This suggests that yielded topics are still not predominantly centred around ESG themes.
+Finally, most of the tested baseline methods score modestly for ESG action orientation, Naiara noticeably higher.
+Yet, Naiara is only able to derive $1$ unique ESG-term, which limits the significance of its analysis.
+These results suggest that the baseline methods may not fully capture information that is material to stakeholders, particularly in terms of conveying ESG actions and efforts.
+2.4 ESGSenticNet Human Evaluation
+In this section, we evaluate ESGSenticNet’s effectiveness in categorising concepts under the correct topics.
+This ensures that during topic analysis, the topic terms matched from ESGSenticNet reliably correspond to their intended ESG topics.
+To assess this, human annotators determine the accuracy of relations between concepts and topics.
+500 concepts are randomly selected from our knowledge base.
+A concept is found within a knowledge triple, possessing a relation with a topic in the format of (concept, relation, ESG topic).
+A correct relation represents that the concept is accurately classified under the ESG topic.
+The accuracy of the relations are evaluated separately for each topic type—pillar, broad, cross-broad, sub, cross-sub, with each of them containing different ESG topics as described in Table 7.
+Human annotators adhere to the following instructions (B provides more details on our annotators and annotation scheme):
+The table below presents various concepts along with their polarities and respective relations toward ESG topics.
+Each concept may be associated with multiple relations across different ESG topics.
+These topics are classified into five topic types: pillar, broad, sub, cross-broad, and cross-sub.
+If a concept has no relevant association with a topic, it is marked as ‘not applicable’.
+Your task is to evaluate whether each concept is correctly related to its corresponding topics.
+For each topic, determine if the relation is the most appropriate compared to the other topics within the same topic type.
+Mark ‘TRUE’ if the relation and topic are correct, and ‘FALSE’ if they are incorrect.
+Additionally, assess the accuracy of the polarity of each concept, and mark ‘TRUE’ if it is correct and ‘FALSE’ if it is incorrect.
+Topic Type Accuracy
+Pillar 95.0%
+Broad 89.3%
+Cross-Broad 86.1%
+Sub 87.8%
+Cross-Sub 90.1%
+Table 4: Accuracy of ESGSenticNet for the relations across all topic types.
+Results from the human evaluation study highlight that ESGSenticNet retains a high degree of accuracy at 86+% for relations across all topic types, as shown in Table 4.
+These findings suggest that ESGSenticNet has reasonable accuracy in categorising concepts under the correct topics.
+Consequently, the topic terms matched from ESGSenticNet during topic analysis corresponds to its related ESG topic with considerable reliability.
+3 Discussion
+3.1 Case Study - Insights offered by ESGSenticNet
+Figure 4: Frequency of ESGSenticNet topics from the flexible run
+Figure 5: Top ESGSenticNet concepts from the flexible run, in support of the sub topics (dark blue), and their corresponding broad topic (light blue)
+The topic analysis conducted by ESGSenticNet in Section 2 is visualised with the Figure 4 & Figure 5, which represent different ways to analyse the text corpora.
+Figure 4 provides a higher level breakdown of the topics present in the corpora.
+It highlights the top 5 frequently occurring broad ESG topics present in the corpora, as well as the proportion of their constituent sub topics, in accordance with ESGSenticNet’s hierarchical taxonomy.
+The topic frequencies are denoted by the occurrence of matched concepts that ‘support’ a topic, leveraging ESGSenticNet’s predefined relations.
+Separately, Figure 5 provides a more granular analysis of the corpora, highlighting matched concepts that explicitly convey ESG actions, thereby communicating important sustainability information.
+Figure 5 displays the top occurring concepts that support the topics shown in Figure 4.
+Put together, Figure 4 and Figure 5 highlight different methods to analyse a sustainability corpus, which can be utilised depending on a stakeholder’s interest and preference.
+While the case study presents insights derived from a test corpus of 319 SGX sustainability reports, the same form of topical analysis can be applied across different corpora of sustainability reports in various real-world scenarios.
+We detail examples of these in the following (to qualify, real-world use cases may extend beyond those listed):
+  • Analysts for ESG funds can leverage ESGSenticNet’s topical insights to evaluate the sustainability of corporations, which is useful for sustainable portfolio construction [35].
+In this context, where timely, transparent, and interpretable assessments are essential, using an on-demand and open-source tool like ESGSenticNet for rule-based topic analysis is advantageous over relying on ESG scores, which are often obscure, inconsistent, and infrequently updated [6, 32].
+  • Regulators can utilise ESGSenticNet to perform sector-wide audits and benchmark companies’ sustainability disclosures against each other.
+For instance, by applying ESGSenticNet across disclosures from firms within a particular industry, regulators can quickly identify under-reporting on material ESG issues (e.g., labour safety in the construction sector, emissions control in the logistics sector) [27].
+This capability enables data-driven supervision and targeted follow-ups.
+Moreover, as ESGSenticNet operates as an interpretable, rule-based method, it provides transparency in how disclosures are assessed—crucial for public accountability and trust in regulatory oversight.
+Additionally, compared to the other ESG topic analysis methods (i.e.
+existing dictionaries and topic modeling), ESGSenticNet’s deployment holds significant advantages in that it extracts – 1) More relevant ESG insights, as demonstrated by the greater proportion of ESG related terms (highest ESG-relatedness score in Table 7).
+2) ESG insights with greater materiality for stakeholders (highest ESG-action orientation score in Table 7).
+3) ESG insights that are organised according to established ESG frameworks (i.e.
+the extracted topic terms from ESGSenticNet encompass ESG topics that are structured according to established ESG themes (e.g., compliance, management), as shown in Figure 12, due to ESGSenticNet’s taxonomy centric organisation (Section 3).).
+These advantages position ESGSenticNet as a topic analysis tool that could potentially be more effective than existing solutions for the aforementioned scenarios.
+3.2 Implications
+Our results indicate that ESGSenticNet outperforms existing methods in capturing a significantly higher proportion of ESG-related and ESG-action oriented topic terms while also identifying a large number of unique ESG terms.
+This suggests that stakeholders using ESGSenticNet for ESG topic analysis can extract more comprehensive and relevant insights from sustainability reports than existing approaches, thereby enhancing ESG assessments.
+Crucially, this capability could enable new discoveries in sustainable finance and sustainability compliance research, such as examining how textual patterns in sustainability reports relate to firm value [71] or sustainability violations [57].
+ESGSenticNet facilitates structured information retrieval by identifying ESG topics—organised according to a hierarchical taxonomy—within a report and surfacing the specific concepts that constitute them.
+This structured output can also support document classification by enabling stakeholders to flag disclosures with distinctive reporting patterns, such as under-reporting in key sustainability areas [27], as discussed in Section 3.1.
+These capabilities collectively support real-world decision-making use cases, including portfolio construction for ESG fund managers [35] and sector-level audits for regulators [27], as outlined in Section 3.1.
+Moreover, our findings highlight that ESGSenticNet can function effectively as a structured lexicon without requiring additional training or tuning, while producing superior results over expensive state-of-the-art NLP topic models.
+This can make ESG topic analysis more accessible to stakeholders with limited technical expertise or computational resources.
+Given the interdisciplinary nature of sustainability research, this accessibility is essential for enabling broader adoption by researchers across different fields, thereby facilitating diverse perspectives and approaches that can drive further discoveries in sustainability research.
+Additionally, ESGSenticNet’s superior performance over existing methods highlights the effectiveness of our scalable and automated neurosymbolic AI approach for lexicon construction in ESG topic analysis.
+Given its effectiveness, this framework could also be adapted to expand lexicons in other key sustainability contexts that require topic analysis—such as climate risk disclosures [49] and environmental policy analysis [12]—where large-scale lexicon development is constrained by high human annotation costs, making automation a necessity.
+Moreover, this scalable and automated framework also allows the knowledge base to be continuously updated at minimal cost, ensuring that the lexicons remain relevant over time.
+This is particularly important in the ESG domain, where sustainability reporting content is dynamic and constantly evolving [22].
+As sustainability reports change, maintaining up-to-date lexicons is essential for accurately capturing emerging ESG themes and patterns.
+4 Methods
+In this section, we introduce the data utilised in this study (Section 4.1), before defining how ESGSenticNet’s triples are organised, laying the groundwork for their subsequent derivation in Section 4.5.
+As explained earlier in the introduction (Section 1), ESGSenticNet comprises knowledge triples in the format of (concept, relation, ESG topic).
+The ESG topics within these triples are drawn from our hierarchical taxonomy in Section 4.3, whose construction methodology is detailed in Section 4.2.
+Additionally, we define the relations between concepts and ESG topics within the triples, along with the rules governing these relations, in Section 4.4.
+4.1 Data
+This study utilises a total of 1,998 sustainability reports from 589 Singapore Exchange (SGX) based companies, in the period of 2015 to 2023 inclusive.
+We randomly divide all the reports into the construction set (1,679 reports) and the evaluation set (319 reports) with a ratio of 85:15, without any overlap between the two sets.
+The construction set is utilised for developing ESGSenticNet, enabling taxonomy derivation (Section 4.2) and knowledge base construction (Section 4.5).
+The evaluation set is reserved for evaluating ESGSenticNet (Section 2).
+All reports are publicly accessible online, and are downloaded in PDF format (via the respective company and company-affiliated websites).
+Thereafter, we processed the report into text format via the PyPDF2 tool.
+For reproducibility, we provide the full details on the reports in the supplementary material and instructions on how to access these reports.
+Specifically, we provide CSV files that specify the companies and the corresponding years of each report.
+Using these details (company name & report year), all sustainability reports can be easily searched for and downloaded online.
+Due to space constraints, we can only provide a sample list of the reports we obtained below in Table 5.
+Company Report Date
+ASIAPHOS LIMITED 2019-12
+MEGROUP LIMITED 2020-03
+LHN LIMITED 2020-09
+CHINA REAL ESTATE GRP LIMITED 2018-06
+CITIC ENVIROTECH LIMITED 2018-12
+JAWALA INC. 2019-07
+BUND CENTER INVESTMENT LIMITED 2018-12
+SAMURAI 2K AEROSOL LIMITED 2021-03
+HOR KEW CORPORATION LIMITED 2019-12
+HALCYON AGRI CORPORATION LIMITED 2021-12
+PROPNEX LIMITED 2019-12
+NAM CHEONG LIMITED 2018-12
+HC SURGICAL SPECIALISTS LIMITED 2019-05
+GKE CORPORATION LIMITED 2019-05
+Table 5: Sample list of SGX sustainability reports provided in the supplementary material.
+4.2 Taxonomy Construction Methodology
+ESG taxonomies, particularly those established by regulatory bodies—i.e.
+GRI [43]—provide structured frameworks for organising sustainability reports, enabling companies to communicate their ESG performance effectively.
+However, in practice, firms may not align their reporting content with these frameworks, as the frameworks may not fully capture the complexities of real-world business operations [54].
+To bridge this gap, we augment existing regulatory taxonomies with a data-driven approach that derives additional ESG topics directly from reporting content, involving BERTopic [29] and human validation.
+This approach first involves tokenising each sustainability report into separate sentences, with each sentence transformed into embeddings by Sentence-BERT [62].
+BERTopic is then run on these embeddings, employing the KeyBERT configuration.
+The topic words are then provided to sustainability experts (details of the sustainability experts can be found in B) who refine the topic words into additional topics.
+Examples of topic words and their inferred topics can be found in Table 6:
+Topic keywords Topic
 strengths, staff, training, opportunities, employees Workplace Development
@@
 sponsorship, sponsor, charity, scholarship Community Empowerment
-Table 1: Examples of Categories Inferred from Topic Keywords
-4.2 Resultant Taxonomy
-Table 2 shows the hierarchical taxonomy derived for ESGSenticNet.
-It contains the definitions of the pillars (Environmental, Social, Governance) and topics used in ESGSenticNet.
-Along with the pillars, the broad topics, and cross-broad topics are bolded, with their constituent sub-topics highlighted in bullet points.
-Beside the broad and cross-broad topics, a letter denotes which pillar the topic falls under – (E) Environmental, (S) Social, (G) Governance.
-Category Description
-Environmental Focuses on impact on the environment.
-This includes, but is not limited to, green innovation, green infrastructure, resource stewardship, energy consumption, waste management, pollution reduction, natural resource conservation, and animal welfare.
-Social Focuses on managing relationships with its employees, suppliers, customers, and the communities.
-This includes, but is not limited to, employee diversity, equity, working conditions, customer relations, partnerships and social engagement.
-Governance Focuses on the system of rules, practices, and processes by which a company is directed and controlled.
-This includes, but is not limited to, corporate oversight, corporate transparency, executive compensation, ethical behavior, and shareholder rights.
-Development (E) Drives innovation and growth in sustainability by creating new technologies or infrastructure for sustainability, or supporting these efforts through green financing.
-$\bullet$ Green Technologies Technological systems, products, and processes that assist in reducing environmental impact.
-$\bullet$ Green Infrastructure Environmentally sustainable buildings, infrastructure, energy solutions and urban systems.
-$\bullet$ Green Financing Financial services that support efforts aimed at environmental sustainability.
-Operations (E) Focuses on optimizing existing business activities for the sole purpose of minimizing environmental impact.
-This includes minimising environmental impact by enhancing current practices in corporate operations, supply chains, production processes, or land management, distinct from the development of innovative solutions and technologies.
-$\bullet$ Sustainable Corporate Operations Environmentally responsible practices within corporate offices.
-Examples include but are not limited to promoting sustainability amongst employees, eco-friendly office lighting.
-$\bullet$ Sustainable Supply Chain Logistics and distribution methods or practices that are environmentally responsible, ensuring sustainability in the movement and handling of goods.
-$\bullet$ Sustainable Production Processes Manufacturing processes or practices that are environmentally responsible.
-$\bullet$ Sustainable Land Management Using and managing land in a way that is environmentally responsible, conserves natural resources and protects biodiversity.
-Ecological Conservation (E) Manages the company’s interaction with natural ecosystems, focusing on minimizing harmful impacts and promoting biodiversity and habitat conservation.
-Emissions Control (E) Targets the reduction of various emissions from company operations, with an emphasis on mitigating climate change impacts or protecting air quality.
-$\bullet$ Climate Emissions Manages greenhouse gas emissions that contribute to global warming, such as CO2 and methane.
-The phrase must explicitly relate to greenhouse gases or related terms such as ’carbon’, ’methane’.
-$\bullet$ Air Quality Emissions Manages emissions of non-greenhouse gases, like toxic gas, particulates and sulfur dioxide, which are critical for local air quality and public health.
-Waste Management (E) Concentrates on managing waste after it has been produced, including the reduction, handling, disposal, and treatment of solid waste and wastewater.
-It involves but is not limited to recycling, reusing materials, and employing wastewater treatment processes to mitigate environmental degradation and facilitate resource recovery.
-$\bullet$ Wastewater Management Focuses on the treatment, reuse and recycling of wastewater or liquid waste to prevent pollution or conserve water resources.
-$\bullet$ Solid Waste Management Addresses the management of solid wastes, promoting reduction, recycling treatment, and responsible disposal to minimize environmental impact and encourage circular economy principles.
-Resource Optimisation (E) Prioritizes the proactive, efficient, and sustainable selection and use of energy, water, and materials.
-This approach focuses on minimizing environmental impact through the strategic choice and utilization of resources.
-Distinct from waste management as it does not primarily deal with waste but rather prevents waste generation by optimizing resource use from the start.
-$\bullet$ Renewable and Efficient Energy Focuses on using energy efficiently, reducing energy consumption and increasing the use of renewable energy sources such as solar, wind, and hydroelectric power to minimize environmental impact.
-$\bullet$ Material Sustainability Prioritizes the use of sustainable and renewable materials, focusing on maximizing their efficiency from the outset.
-Distinct from recycling and reuse, which are part of waste management, this approach aims to reduce environmental impact before materials enter the waste stream.
-$\bullet$ Water Conservation Focuses on using water efficienctly, reducing water usage and enhancing sustainable water management practices to minimize initial consumption.
-Distinct from water recycling and reusing, which address management of wastewater.
-Workplace (S) Focuses on the social aspects of sustainability within the company to promote equity, inclusivity, or quality of life, emphasizing employee well-being, inclusivity, or professional growth.
-$\bullet$ Workplace Wellness Efforts that focus on the satisfaction, emotional and physical well-being of employees.
-$\bullet$ Workplace Diversity Efforts that emphasise gender and racial diversity and fairness in the workplace, such as anti-discrimination measures, fair hiring practices etc.
-$\bullet$ Workplace Development Efforts that focus on the skill training and professional development of staff.
-Communications (E/S/G) Focuses on how a company transparently and honestly communicates and discloses its activities.
-Outreach (S) Enhances social sustainability by contributing to community well-being, fostering customer satisfaction, building customer relationships, or collaborating with strategic partners to promote equity, inclusivity, and quality of life.
-$\bullet$ Community Empowerment Contributions to the welfare of local communities through various forms of support and engagement.
-$\bullet$ Strategic Partnerships Collaborations with stakeholders like NGOs, businesses, and charities to achieve shared sustainability goals.
-$\bullet$ Customer Engagement Fostering customer satisfaction, strong relationships and reputation with customers.
-Management (G) Ensures strategic, ethical, or effective governance through rigorous oversight, diverse board composition, active stakeholder engagement, or high ethical standards within company leadership.
-$\bullet$ Management Composition Structures the board or management with diverse expertise and backgrounds to strengthen decision-making and governance effectiveness.
-$\bullet$ Management Ethics Enforces strict ethical standards within company leadership, ensuring management actions uphold corporate integrity.
-$\bullet$ Management Compensation Aligns management, executive and board remuneration with company performance, sustainability or ethical objectives.
-$\bullet$ Stakeholder Engagement Engages key stakeholders to align board or management decisions with broader interests and feedback.
-$\bullet$ Oversight Involves the systematic review and evaluation of management actions to ensure they align with the company’s performance goals, sustainability goals and ethical standards.
-Compliance (G) Ensures adherence to laws, regulations, or standards across environmental protection, labor rights, consumer safety, ethical conduct, financial practices, or data protection, distinguishing itself from voluntary sustainability initiatives by focusing on mandatory requirements.
-$\bullet$ Environmental Compliance Adherence to environmental laws such as for emissions, water pollution, protection of biodiversity etc.
-$\bullet$ Worker & Consumer Safety Adherence to laws that guarantee the safety of products and services to consumer health, as well as fair labor practices and the protection of workers’ rights.
-$\bullet$ Business Compliance Adherence to anti-corruption, anti-competitive practices, financial regulations, or financial reporting.
-$\bullet$ Data Privacy & Cybersecurity Protection Protection and confidentiality of data, which includes measures against cyber threats and compliance with data security regulations.
-Table 2: ESGSenticNet Sustainability Taxonomy
-5 ESGSenticNet
-ESGSenticNet comprises knowledge triplets with the format of (concept, relation, category), following from [27].
-To form the knowledge triplets, concepts within our knowledge base are labelled according to their relations (table 3) with the categories in our taxonomy (table 2).
-The relations, ‘supports’ and ‘undermines’, clarify how concepts advance or impede crucial aspects of corporate sustainability.
-In contrast, the relation ‘aligns with’ does not consider a concept’s impact, highlighting a concept’s general connection with broader pillars of environmental, social or governance.
-Additionally, categories are divided according to their different category types–pillar, broad, cross-broad, sub, cross-sub, as shown in table 4, with relations assigned according to rules (section 5.1).
-ESGSenticNet is constructed through a pipeline that involves three main phases: parsing concepts from the text corpus (section 5.2), processing parsed concepts to ensure coherence (section  5.3), labelling processed concepts according to their relation with respect to our developed taxonomy (sections 5.55.65.45.7).
-Concept parsing involves extracting candidate concepts from sustainability disclosures through dependency parsing and part-of-speech tagging.
-From these candidates, coherent concepts are filtered through text parsing followed by leveraging GPT-4 [3] (section 5.2).
-The filtered concepts undergo a labelling methodology that we have developed to minimise computational cost while enhancing the diversity of labelled concepts.
-Accordingly, this labelling methodology (figure 1 1 through 4) entails leveraging semantic graphs, clustering techniques, seed selection algorithms, and label propagation.
-This methodology allows us to label a large amount of diverse concepts despite performing inference on a subset of the concept samples.
-As a result, the extensiveness and diversity of our knowledge base is significantly increased, which is imperative to its application for corporate sustainability analysis.
+Table 6: Examples of Topics Inferred from Topic Keywords
+Additional topics such as these are then added to an adapted version of the existing SASB framework [25], forming a taxonomy of ESG topics (Section 4.3) utilised in ESGSenticNet.
+4.3 Resulting Taxonomy
+Topic Type Topics
+Pillar ‘Environmental’, ‘Social’, ‘Governance’.
+Broad ‘Development’, ‘Operations’, ‘Workplace’, ‘Outreach’, ‘Management’, ‘Communications’, ‘Compliance’
+Sub ‘Green Technologies’, ‘Green Infrastructure’, ‘Green Financing’, ‘Sustainable Corporate Operations’, ‘Sustainable Supply Chain’, ‘Sustainable Production Processes’, ‘Sustainable Land Management’, ‘Workplace Wellness’, ‘Workplace Diversity’, ‘Workplace Development’, ‘Community Empowerment’, ‘Strategic Partnerships’, ‘Customer Engagement’, ‘Management Composition’, ‘Management Ethics’, ‘Management Compensation’, ‘Stakeholder Engagement’, ‘Oversight’, ‘Environmental Compliance’, ‘Worker & Consumer Safety’, ‘Business Compliance’, ‘Data Privacy & Cybersecurity Protection’
+Cross-Broad ‘Resource Optimisation’, ‘Waste Management’, ‘Emissions Control’, ‘Ecological Conservation’
+Cross-Sub ‘Renewable and Efficient Energy’, ‘Material Sustainability’, ‘Water Conservation’, ‘Wastewater Management’, ‘Solid Waste Management’, ‘Climate Emissions’, ‘Air Quality Emissions’
+Table 7: Topic types and topics, with the different topics delineated by a comma
+We summarise the hierarchical taxonomy derived for ESGSenticNet in Table 7.
+The taxonomy comprises of topics are divided according to their different topic types–pillar, broad, cross-broad, sub, cross-sub, as shown in Table 7.
+The taxonomy is structured hierarchically, with topics categorised into different levels based on their scope and relationship.
+At the highest level, all topics fall under the broad ESG pillars—Environmental (E), Social (S), and Governance (G)—which represent the overarching dimensions of sustainability.
+Beneath the pillars, topics are grouped into broad and cross-broad topics, which define major sustainability themes.
+Each broad or cross-broad topic may have constituent sub-topics, which represent more specific areas within that theme.
+While Table 7 summarises our taxonomy, it does not detail the constituent sub-topics of each broad or cross-broad topic.
+For these relationships as well as topic definitions, please refer to the full descriptions in A and Table 16.
+We elaborate more on the features of this taxonomy below:
+  • Realistic: Given that we leverage the empirical content from the sustainability reports, our taxonomy comprises additional ESG topics that might otherwise be overlooked by standard frameworks.
+For instance, ‘Green Technology’ is derived as a taxonomy topic, while not being explicitly emphasised in standard ESG reporting frameworks—i.e.
+GRI [43], SASB [25].
+  • Activities Cross-Theme: To capture the interconnectedness of sustainability dimensions [48], additional topics include areas of a firm’s business activities (‘operations’, ‘development’ etc.).
+In contrast with existing frameworks that primarily describe direct sustainability impacts (‘emissions control’, ‘waste management’ etc.), this offers insight into the sustainability effects of specific business activities.
+In our taxonomy, topics pertaining to strategic business activities fall under ‘broad’ & ‘sub’ topic types, and are kept separate from topics pertaining to direct sustainability impacts, which fall under ‘cross-broad’ & ‘cross-sub’ topic types.
 4.4 ESGSenticNet Relations & Rules
+With the taxonomy established (Table 7), we now define the relation types between concept and pillar/topic (Table 8), as well as the rules governing these relations.
+Each concept within a triple can possess one of the following relations with a pillar/ topic:
 Relation Description
 aligns with The concept is positively or negatively related to a specific Environmental, Social, or Governance (ESG) pillar.
-supports The concept directly advances the sustainability topic with a clear and immediate positive impact.
-undermines The concept directly impedes the sustainability topic with a clear and immediate negative impact.
+supports The concept directly advances the ESG topic with a clear and immediate positive impact.
+undermines The concept directly impedes the ESG topic with a clear and immediate negative impact.
 Table 8: Relations Types
-Category Type Categories
-Pillar ‘Environmental’, ‘Social’, ‘Governance’.
-Broad ‘Development’, ‘Operations’, ‘Workplace’, ‘Outreach’, ‘Management’, ‘Communications’, ‘Compliance’
-Sub Children categories of broad categories
-Cross-Broad ‘Resource Optimisation’, ‘Waste Management’, ‘Emissions Control’, ‘Ecological Conservation’
-Cross-Sub Children categories of cross-broad categories
-Table 4: Categories and category types
-ESGSenticNet is defined with the following relations (table 3), category types (table 4), as well as the following rules.
+The relations, ‘supports’ and ‘undermines’, clarify how concepts advance or impede crucial topics of sustainability.
+In contrast, the relation ‘aligns with’ does not consider a concept’s impact, highlighting a concept’s general connection with broader pillars of environmental, social or governance.
+At this juncture, it is worthwhile to qualify that the deployment of ESGSenticNet during topic analysis solely uses the ‘supports’ relation (explained further in Section 2.2), while the other relations serve as supplementary information.
+Separately, we now outline the rules that govern all relations in ESGSenticNet:
   • Pillar Assignment: Each concept ‘aligns with’ a specific pillar, mapping the concepts to the broader sustainability context.
-Single Label within Category Types: Each concept can hold only one relation with one category within each category type.
-  • Cross-Labels between Parent & Children: Each concept can hold relations with higher-level categories (‘broad’ and ‘cross-broad’), as well as with more detailed categories (‘sub’ and ‘cross-sub’).
-  • Cross-Labels between Cross & Non-Cross: Concepts that ‘align with’ the ‘environmental’ pillar can hold relations with ‘cross-broad’ and ‘cross-sub’ categories, in addition to ‘broad’ and ‘sub’ categories.
-This reflects the multi-faceted nature of environmental concepts, which can simultaneously impact environmental stewardship (represented by ‘cross-broad’, ‘cross-sub’ categories) and strategic business activities (represented by ‘broad’ and ‘sub’ categories).
+  • Single Label within Topic Types: Each concept can hold only one relation with one topic within each topic type.
+  • Cross-Labels between Parent & Children: Each concept can hold relations with higher-level topics (‘broad’ and ‘cross-broad’), as well as with more detailed topics (‘sub’ and ‘cross-sub’).
+  • Cross-Labels between Cross & Non-Cross: Concepts that ‘align with’ the ‘environmental’ pillar can hold relations with ‘cross-broad’ and ‘cross-sub’ topics, in addition to ‘broad’ and ‘sub’ topics.
+This reflects the multi-faceted nature of environmental concepts, which can simultaneously impact environmental stewardship (represented by ‘cross-broad’, ‘cross-sub’ topics) and strategic business activities (represented by ‘broad’ and ‘sub’ topics).
+4.5 ESGSenticNet Construction through a Neurosymbolic Framework
+In sections 4.1 4.2 4.3 4.4, we outlined the preliminaries by introducing the data and defining the ESG topics and relations within the knowledge triples (concept, relation, ESG topic) in ESGSenticNet, thereby establishing ESGSenticNet’s organisational structure.
+In this section, we go beyond definitions to detail our neurosymbolic framework for deriving the triples, explaining how concepts are identified and assigned their respective relations.
+Figure 6: Overview of neurosymbolic framework for ESGSenticNet construction
+The neurosymbolic framework for constructing ESGSenticNet involves three main phases as shown in figure 6.
+Step 1: ESG Concept Parser (Section 4.6), involves concept parsing, which encompasses extracting candidate concepts from sustainability reports through dependency parsing and part-of-speech tagging.
+Step 2: Quality Control (Section 4.7), involves processing coherent concepts from these candidate concepts to ensure their coherence, as a form of quality control.
+Step 3: Labelling Process (Section 4.8), involves a labelling methodology to classify the filtered concepts according to the topics within our constructed taxonomy in Section 4.3, leveraging the rules defined in Section 4.4.
+Accordingly, this labelling methodology in step (3) is detailed in Figure 6 (following processes i through iv).
+These processes integrate GPT-4o inference with a graph-based semi-supervised method, as explained below:
+  1. (i)
+    Semantic Graph Construction (Section 4.8.1): A semantic graph is constructed using the filtered concepts from step (2), leveraging their semantic similarities.
+  2. (ii)
+    Seed Selection (Section 4.8.2): Seed concepts are selected from the constructed semantic graph using a seed selection algorithm that applies clustering techniques and graph analysis.
+  3. (iii)
+    Seed Annotation (Section 4.8.3): Using GPT-4o, the relations between selected seed concepts and taxonomy topics are labelled.
+  4. (iv)
+    Label Propagation (Section 4.8.4): The relation labels from seed concepts are propagated to other concepts through a label propagation algorithm.
+Upon completing the labelling process, ESGSenticNet is fully constructed.
+Notably, this labelling process allows us to label a large amount of diverse concepts despite performing inference on a subset of the concept samples.
+As a result, computational cost constraints are mitigated, as we are able to label concepts more extensively.
+This, in turn, significantly increases the extensiveness and diversity of our knowledge base, enhancing its effectiveness as a lexical method for ESG topic analysis.
 4.6 ESG Concept Parser
@@
 In the context of sustainability analysis, multiword expressions frequently communicate key sustainability ideas [65].
-We refer to these multiword expressions as sustainability ”concepts”.
-Extracting these concepts or ”concept extraction” has been explored for metaphors [36], finance [18] and sentiment analysis [10].
-Yet, to the best of our knowledge, concept extraction within the sustainability domain has not been studied.
-Therefore, to capture these sustainability concepts, a new concept parser is required, with an explicit focus on the distinct elements of sustainability discourse.
+We refer to these multiword expressions as ESG “concepts”.
+Extracting these concepts or “concept extraction” has been explored for metaphors [46], finance [24] and sentiment analysis [13].
+Yet, to the best of our knowledge, concept extraction within the ESG domain has not been studied.
+Therefore, to capture ESG concepts, a new concept parser is required, with an explicit focus on the distinct elements of sustainability discourse.
 These include the following designs:
@@
 verbs of volition like ‘plan’, ‘commit’) is observed when a company pledges future sustainability activities [64].
-To effectively extract these unique aspects of sustainability language, we formalise sustainability concepts, or ESG concepts, through the following syntactic structure–verbs acting on complex noun phrases, where the words involved function as modifiers, descriptive and prescriptive elements.
-Specifically, the concepts are two or three word phrases, typically beginning with a $verb$, followed by $noun+noun$, or $adjective+noun$ combinations with specific grammatical dependencies.
-These dependencies include: $verb-noun$ connection via (nsubj, obj, obl), $adjective-noun$ or $noun-noun$ connection via (compound, amod, nn, appos, flat, nmod).
-This framework underpins the ESG Concept Parser algorithm (1), designed to capture key ideas within sustainability disclosures, particularly sustainability efforts.
-For instance, ‘improve (VB) workplace (NN) safety (NN)’, where ‘workplace’ is the $compound$ of ‘safety’ and ‘safety’ is the $dobj$ of ‘improve’.
+To effectively extract these unique aspects of sustainability language, we formalise ESG concepts through the following syntactic structure–verbs acting on complex noun phrases, where the words involved function as modifiers, descriptive and prescriptive elements.
+Specifically, the concepts are two or three word phrases, typically beginning with a verbverb, followed by $noun+noun$, or $adjective+noun$ combinations with specific grammatical dependencies.
+These dependencies include: verbnounverb-noun connection via (nsubj, obj, obl), adjectivenounadjective-noun or nounnounnoun-noun connection via (compound, amod, nn, appos, flat, nmod).
+This framework underpins the ESG Concept Parser algorithm (1), designed to capture key ideas within sustainability reports, particularly ESG efforts.
+For instance, ‘improve (VB) workplace (NN) safety (NN)’, where ‘workplace’ is the compoundcompound of ‘safety’ and ‘safety’ is the dobjdobj of ‘improve’.
 Algorithm 1 ESG Concept Parser
-1:Input: Sentence from sustainability disclosure
+1:Input: Sentence from sustainability report
 2:Output: ESG Concepts
@@
 15:     if dependent is noun then
-16:                 add to Phrases: $word_{1}$+dependent+$word_{2}$ as verb+noun+noun combination
+[…]
 17:     end if
 18:     if dependent is adj then
-19:                 add to Phrases: $word_{1}$+dependent+$word_{2}$ as verb+adj+noun combination
+[…]
 20:     end if
@@
 23:end for
-5.3 Processing concepts for coherence
-To minimise computational cost, parsed ESG concepts from algorithm (1) are filtered to retain the top 110k most frequently occurring ones.
+4.7 Quality Control
+To minimise computational cost, parsed concepts from algorithm (1) are filtered to retain the top 110k most frequently occurring ones.
 Of these, we eventually annotate 23k concepts based on our estimated computational resources, which dictates the number of seeds we select within our semantic graph.
-Prior to the annotation process, these concepts must undergo processing to ensure their coherence.
-This involves utilising GPT-4o in few-shot setting (Brown, 2020) to ensure their coherence.
+Prior to the annotation process, these concepts must undergo processing to ensure their coherence, as a form of quality control.
+This involves utilising GPT-4o in few-shot setting (Brown, 2020).
 GPT-4o is leveraged due to the capacity of large generative LLMs to infer [72], as well as their powerful natural language understanding capabilities [76].
-Using GPT-4o to process concepts for coherence involves conducting the following inference tasks in order: (1) reorder words, if required, to improve the coherence of the ESG concept, (2) determine if the subsequent ESG concept is intelligible.
-We provide the prompt for (1) and (2) through figures 3 and 4 respectively.
+Using GPT-4o to process concepts for coherence involves conducting the following inference tasks in order: (1) reorder words, if required, to improve the coherence of the concept, (2) determine if the subsequent concept is intelligible.
+We provide the prompt for (1) and (2) through Figure 7 and Figure 8 respectively.
 [Task Description] As a language expert, your task is to rearrange words in phrases to enhance their coherence and intelligibility if necessary.
@@
 Figure 8: Prompt for evaluating an ESG concept’s coherence
-5.4 Semantic Graph Construction
+4.8 Labelling Process
+4.8.1 Semantic Graph Construction
 $G=(V,E)$ (1)
@@
 $E=\{(\mathbf{e}_{i},\mathbf{e}_{j})\mid\text{sim}(\mathbf{e}_{i},\mathbf{e}_{j})>0.80\}$ (4)
-Processed ESG concepts undergo labelling to derive their relationships with different sustainability topics.
-Given the exceedingly large number of ESG concepts, annotating all concepts would be prohibitively expensive.
-To overcome this, we devise a semi-supervised framework that enables a large volume of diverse ESG concepts to be labelled from annotating only a subset of all ESG concepts.
-Via a semantic graph, we leverage the semantic similarities between ESG concepts to propagate labels from annotated concepts (seeds) to unlabelled concepts (non-seeds).
+$w_{ij}=\text{sim}(\mathbf{e}_{i},\mathbf{e}_{j}),\quad\forall(\mathbf{e}_{i},\mathbf{e}_{j})\in E$ (5)
+Processed concepts undergo labelling to derive their relationships with different ESG topics.
+Given the exceedingly large number of concepts, annotating all concepts would be prohibitively expensive.
+To overcome this, we devise a semi-supervised framework that enables a large volume of diverse concepts to be labelled from annotating only a subset of all concepts.
+Via a semantic graph, we leverage the semantic similarities between concepts to propagate labels from annotated concepts (seeds) to unlabelled concepts (non-seeds).
 This graph is constructed by transforming each concept $cp_{i}$ into embeddings $e_{i}$ through S-BERT [62], with edges constructed by leveraging a cosine similarity threshold of 0.80 between each node $e_{i}$.
-5.5 Seed Selection
+To qualify the strength of relation between nodes (concepts), the edge weights $w_{ij}$ are defined as the cosine similarity scores between the nodes.
+We describe the full graph via equations (1) through (5).
+4.8.2 Seed Selection
 Within our semantic graph, seeds are selected for annotation through data clustering and graph analysis.
 Given the semi-supervised label propagation from seeds to non-seeds, we formulate the following objectives for seed selection: (I) Seeds must be semantically diverse to ensure that diverse labelled samples are yielded post label propagation.
-This enhances the variety and comprehensiveness of labelled ESG concepts in our knowledge base.
+This enhances the variety and comprehensiveness of labelled concepts in our knowledge base.
 (II) Seeds must propagate their labels to as many non-seeds as possible.
 This maximises the number of labelled samples we can attain while minimising annotation costs.
-5.5.1 Data Clustering
+To achieve (I) and (II), we utilise Data Clustering and Graph Analysis, with these components culminating in a Seed Selection Algorithm.
 $CQI=\frac{|\{e\in V\mid\alpha_{e}>0.60\}|}{|V|}$ (6)
-To accomplish (I), the embeddings of ESG concepts $e_{i}$ are clustered to enable seed selection from diverse semantic spaces.
+Data Clustering.
+To accomplish (I), the embeddings of concepts $e_{i}$ are clustered to enable seed selection from diverse semantic spaces.
 This clustering involves projecting the embeddings to a lower dimensional space via UMAP [47], before running HDBSCAN [15] on these projected embeddings.
@@
 In our later section, label propagation leverages the default dimensions of S-BERT embeddings.
-Clusters are evaluated via a Confidence Quality Index ($CQI$), which denotes the proportion of embeddings, $e$, out of all embeddings, $V$, with a clustering confidence level $\alpha_{e}$ beyond 0.60.
+Clusters are evaluated via a Confidence Quality Index ($CQI$), which denotes the proportion of embeddings, $e$, out of all embeddings, $V$, with a clustering confidence level $\alpha_{e}$ beyond 0.60, as described by equation (6).
 For our use of HDBSCAN, this simple metric is more appropriate than others like Slihouette Score [63], which assumes that all embeddings belong to a cluster.
-Hyperparameter tuning is done via random search, with the best run having a $CQI$ of 0.66, UMAP parameters $n$_$neighbours$$=$32, $n$_$components$$=$2, and the HDBSCAN parameter $min$_$cluster$_$size$$=$2.
-5.5.2 Graph Analysis
+Hyperparameter tuning is done via random search, with the best run having a $CQI$ of 0.66, UMAP parameters nn_neighboursneighbours==32, nn_componentscomponents==2, and the HDBSCAN parameter minmin_clustercluster_sizesize==2.
 $\begin{split}Q(\mathbf{e}_{i})=|\left\{(\mathbf{e}_{i},\mathbf{e}_{j})\in E\mid(\mathbf{e}_{i}\in\mathbf{c_{k}})\right.\\ \left.\wedge\left(\forall\mathbf{e}_{s}\in S_{c_{k}},(\mathbf{e}_{s},\mathbf{e}_{j})\notin E\right)\right\}|\end{split}$ (7)
+Graph Analysis.
 To accomplish (II), graph analysis is utilised to select seeds that are the most consequential nodes.
@@
 Unique edges are defined as connections to nodes that do not share an edge with the already selected seed nodes within the same cluster.
-We formulate the local consequential score $Q(e_{i})$, where $S_{c_{k}}$ is the set of nodes already selected as seeds within the same cluster, $c_{k}$, as the node $e_{i}$.
-5.5.3 Seed Selection Algorithm
+We formulate the local consequential score $Q(e_{i})$, where $S_{c_{k}}$ is the set of nodes already selected as seeds within the same cluster, $c_{k}$, as the node $e_{i}$, with this described in equation (7).
+Seed Selection Algorithm.
 Capitalising on Data Clustering and Graph Analysis, an algorithm (2) is developed for seed selection to achieve objectives (I) & (II).
@@
 2:Output: $S_{\text{total}}$
-3:$N\leftarrow$ $|V|$
-4:$P\leftarrow\frac{T}{N}$
-5:$\Delta_{prev}\leftarrow\infty$
+[…]
+[…]
+[…]
 6:while True do
@@
 30:return $S_{\text{total}}$
-5.5.4 Effectiveness of ESGSenticNet Seed Selection Algorithm
-Seed Selection Method Propagated Labels
-Highest Degree Centrality 4686
-Random Selection 7456
-ESGSenticNet 9617
-Table 5: Effectiveness of Label Propagation, based on the number of newly propagated labels
-In this baseline study, we assess the extent of label propagation through ESGSenticNet’s seed selection algorithm, against naive methods like selecting seeds randomly or based on highest degree centrality scores.
-Each method is run on the constructed semantic graph (section 5.4), to select different batches of 10k seeds.
-Aligning with majority of the classification tasks in the main study, three different psuedo-labels are used, and randomly assigned to the selected seeds.
-Label propagation is run to observe the number of newly propagated labels.
-Although propagation would be more accurate with real labels, random pseudo-labelling provides an unbiased estimate of propagation extensiveness irrespective of label content, and reduces the cost of extensive testing.
-Moreover, the entire process is repeated over 100 trials, showing that across each method, the number of newly propagated labels remain consistent despite the randomness of seed labelling.
-Compared to naive methods, we observe that ESGSenticNet’s seed selection algorithm results in a greater number of newly propagated labels.
-5.6 Seed Annotation
+4.8.3 Seed Annotation
 Figure 9: Flowchart depicting inference process
-Through GPT-4o few-shot prompting, selected seeds are annotated with their relations (‘aligns with’, ‘supports’, ‘undermines’) with respect to a pillar or category, before their labels are propagated within the semantic graph.
-Similar to popular sustainability lexicons [6, 54], ESGSenticNet’s approach focuses on seed annotation before leveraging textual similarities to expand the initial seed corpus.
-Yet, while approaches in the literature annotate a small number of seeds, ESGSenticNet annotates an extensive number of seeds ($>16k$) through GPT-4o.
-Consequently, ESGSenticNet retains a significant proportion of seed concepts directly labelled through GPT-4o, while non-seeds that rely on textual similarity for labelling are less prevalent.
-This enriches the diversity of ESGSenticNet, given that the reasoning capabilities of GPT-4o [24] allows it to infer relations for concepts outside of similar semantic spaces.
+Through GPT-4o few-shot prompting, the selected seeds are annotated with their relations (‘aligns with’, ‘supports’, ‘undermines’) with respect to a pillar or topic, before their labels are later propagated within the semantic graph (Section 4.8.4).
+ESGSenticNet retains a considerable number of seed concepts (>16k>16k) directly labelled through GPT-4o.
+This improves the accuracy of ESGSenticNet’s relations, given that the reasoning capabilities of GPT-4o [36] allows it to infer relations for concepts outside of similar semantic spaces.
 For example, consider the concepts ‘use led light’, ‘achieve energy efficiency’ categorised under ‘Resource Optimisation’.
@@
 However, using GPT-4o, ‘emphasise water importance’ is appropriately labelled, due to the LLM’s capacity to infer semantic relationships beyond mere textual similarities.
-Moreover, GPT4o as an automated technique allows for a large number of concepts to be labelled, overcoming prohibitive human labelling costs.
-This allows us to derive an extensive knowledge base.
 To improve GPT-4o classification accuracy, inference tasks are divided into smaller, more manageable sub-tasks [17].
 These tasks align with the taxonomy’s hierarchy levels, following the flowchart in Figure 9.
-Subsequent categorisations of seeds are contingent upon their initial classification into broader categories, and it aligns with the predefined associations between the ESG pillars and broad categories, as well as the connections between the broad and sub categories.
+Subsequent categorisations of seeds are contingent upon their initial classification into broader topics, and it aligns with the predefined associations between the ESG pillars and broad topics, as well as the connections between the broad and sub topics.
 Prompts for these tasks are highlighted in Figure 10 (pillar labelling), Figure 11 (relation-topic labelling).
 [Task Description]
-As a sustainability expert, your task is to classify phrases into the categories of Environmental, Social, Governance, or Others based on their explicit relevance to the Environmental, Social, Governance topics.
-[Definitions of Categories]
+As a sustainability expert, your task is to classify phrases into the topics of Environmental, Social, Governance, or Others based on their explicit relevance to the Environmental, Social, Governance topics.
+[Definitions of topics]
 Definitions of ESG according to taxonomy, with the inclusion of ‘Others’
 Others: Reserved for phrases that do not meet the explicitness and specificity requirements for Environmental, Social, or Governance.
-This includes phrases that are vague, ambiguous, or cover multiple Environmental, Social, Governance topics without clear categorization.
+This includes phrases that are vague, ambiguous, or cover multiple Environmental, Social, Governance topics without clear categorisation.
 [Response Instructions]
-Evaluate each phrase’s explicit content to determine its most appropriate category.
-Assign each phrase to only one category: Environmental, Social, Governance, or Others.
-Only phrases that are clear and explicit can be categorized under Environmental, Social, or Governance.
-Phrases that are vague or not clearly related to these topics should be categorized as Others.
+Evaluate each phrase’s explicit content to determine its most appropriate topic.
+Assign each phrase to only one topic: Environmental, Social, Governance, or Others.
+Only phrases that are clear and explicit can be categorised under Environmental, Social, or Governance.
+Phrases that are vague or not clearly related to these topics should be categorised as Others.
 Ensure all responses strictly adhere to the definitions and format provided.
@@
 Figure 10: Prompt for labelling an ESG concept’s pillar
-[Task Description] As a sustainability expert, you are tasked with evaluating the relationship between phrases and sustainability-related topics.
+[Task Description] As a sustainability expert, you are tasked with evaluating the relationship between phrases and ESG-related topics.
 [Definitions of Relationships] supports: The phrase, specific and explicit to the topic, directly advances it with a clear and immediate positive impact.
@@
 Figure 11: Prompt for labeling an ESG concept’s relation with a topic
-5.7 Label Propagation Algorithm
+4.8.4 Label Propagation
 $L^{(k+1)}=D^{-1/2}AD^{-1/2}L^{(k)},$ (8)
@@
 This increases the labelled data within our knowledge base without incurring additional labelling costs.
+We utilise the label propagation algorithm for its ability to perform soft transducive inference — iteratively proliferating labels across the graph probabilistically, based on global similarity [69].
+This contrasts with other algorithms — i.e.
+k-nearest neighbours [59], which assign hard labels without modeling uncertainty (i.e.
+via probability).
+In our context, the label propagation algorithm’s incorporation of probability allows it to better capture semantic nuances and ambiguity across concepts.
+This is useful when one concept may relate to multiple categories.
 Specifically, the following algorithm is iterated until convergence, where $A$ is an adjacency matrix containing $a_{i,j}$ that denotes the edge weight between connected nodes $e_{i}$ and $e_{j}$.
@@
 $m$ is the number of nodes possessing labels and $D$ is a diagonal matrix containing $d_{i,i}$, which equals the sum of elements in row $i$ of $A$.
-We use parameters $n\_layers$=50, $\alpha$=0.5.
-5.8 ESGSenticNet Statistics
-Categories Total Triplets Number of triplets with the relations
+We use parameters n_layersn\_layers=50, $\alpha$=0.5.
+4.9 ESGSenticNet — Statistics & Samples
+Figure 12: Overview of ESGSenticNet
+After label propagation, ESGSenticNet is fully constructed.
+We provide an visualisation of ESGSenticNet in Figure 12.
+ESGSenticNet comprises 44232 triples containing 23245 unique ESG concepts.
+Of the 23245 concepts, 16011 are seeds and 7234 are non-seeds.
+We provide samples of ESGSenticNet’s knowledge triples (concept, relation, ESG topic) in Table 9.
+Concept Relation ESG topic
+build diverse workplace supports workplace diversity
+halve carbon emission supports emissions control
+involve workplace injury undermines worker & consumer safety compliance
+minimise resource consumption supports resource optimisation
+organise charity event supports outreach
+produce chemical waste undermines operations
+Table 9: Samples of knowledge triples from ESGSenticNet
+Table 10 provides a meta statistic of of how ESGSenticNet compares with existing sustainability dictionaries—i.e Baier [5], Naiara [60], Kang [39].
+To qualify, the concepts within ESGSenticNet’s triples are taken as “terms”.
+Sustainability Dictionaries/ Knowledge Base(s) Total Unique Terms
+ESGSenticNet 23245 (within 44232 triples)
+Baier 491
+Naiara 167
+Kang* 792
+Table 10: Total number of terms of different sustainability corpora, *indicates not publicly available.
+Finally, we provide a full statistical breakdown of ESGSenticNet in Appendix C Table 17. This outlines the total triples corresponding to each ESG topic, with a breakdown of the relations involved.
+4.10 Ablation Studies
+4.10.1 Effectiveness of ESGSenticNet Seed Selection Algorithm
+In this baseline study, we assess the extent of label propagation through ESGSenticNet’s seed selection algorithm, against naive methods like selecting seeds randomly or based on highest degree centrality scores.
+Each method is run on the constructed semantic graph (Section 4.8.1), to select different batches of 10k seeds.
+Aligning with majority of the classification tasks in the main study, three different psuedo-labels are used, and randomly assigned to the selected seeds.
+Label propagation is run to observe the number of newly propagated labels.
+Although propagation would be more accurate with real labels, random pseudo-labelling provides an unbiased estimate of propagation extensiveness irrespective of label content, and reduces the cost of extensive testing.
+Moreover, the entire process is repeated over 100 trials, showing that across each method, the number of newly propagated labels remain consistent despite the randomness of seed labelling.
+Seed Selection Method Propagated Labels
+Highest Degree Centrality 4686
+Random Selection 7456
+ESGSenticNet 9617
+Table 11: Effectiveness of Label Propagation, based on the number of newly propagated labels
+Compared to naive methods, from Table 11, we observe that ESGSenticNet’s seed selection algorithm results in a greater number of newly propagated labels.
+4.10.2 GPT-4 Human Expert Agreement on Topic Analysis Evaluation Metrics
+GPT-4 evaluations have shown significant reliability in the literature [80].
+In line with established research practices [61], we validate the GPT-4 evaluations in our use case by conducting a GPT4-human agreement study.
+Specifically, a study is conducted to assess GPT-4’s ability for the following tasks– i) classifying topic terms as ESG-related, ii) classifying topic terms as actions taken toward ESG.
+A random sample of 510 topic terms were gathered for each task.
+GPT-4 annotations for each set of topic terms are compared with annotations from 2 different groups of 3 human experts each, with each group handling a different task.
+These human experts comprise of doctoral candidates actively conducting research in sustainability finance, and thus possess domain expertise well beyond the level required for this validation task.
+Moreover, to ensure the robustness of their judgements, the human experts were deliberately kept independent from the paper’s development.
+An average agreement score of 83.7% observed for task i) and 91.8% for task ii).
+While this highlights a reasonable level of robustness for GPT-4 evaluations in our study, it also underscores the broader potential of GPT-4 as a scalable and reproducible evaluation tool in the ESG domain.
+For evaluating GPT-4 human agreement, the following instructions were given for classifying topic terms as ESG-related.
+The table below presents various terms. Determine if each term is ESG-related. Mark ‘TRUE’ if the term is ESG-related and ‘FALSE’ otherwise.
+For evaluating GPT-4 human agreement, the following instructions were given for classifying topic terms as an action taken toward ESG.
+The table below presents various terms.
+Determine if each term expresses an action toward improving a company’s ESG performance.
+Mark ‘TRUE’ if the term express an action toward improving ESG performance and ‘FALSE’ otherwise.
+4.10.3 Effectiveness of ESGSenticNet Flexible Matching Method
+A study is conducted to assess the precision of matching ESGSenticNet concepts through the ‘flexible’ method depicted in Section 2.2.
+A random sample of 200 sentences matched through this method was collected, excluding sentences matched via the ‘precise’ method.
+Human annotators were tasked with evaluating whether each sentence clearly conveys the idea expressed represented by its matched concept.
+From our results, 75.0% of sentences comprised the idea expressed by its corresponding matched concepts.
+This suggests that the ‘flexible’ method retains a moderately high level of precision, despite allowing for greater semantic variability.
+For evaluating the ESGSenticNet flexible matching method, the following instructions were given to evaluate if each sentence clearly conveys the idea expressed by its matched concept.
+The table below presents different sentences and a corresponding concept.
+Determine if each sentence clearly conveys the idea expressed by its matched concept.
+Mark ‘TRUE’ if the sentence clearly conveys the concept’s idea, and ‘FALSE’ otherwise.
+5 Literature Review
+5.1 NLP-driven ESG topic analysis.
+To extract insights from sustainability reports, NLP tools can be deployed for different tasks, such as ESG topic analysis [28, 41], sentiment classification [58], extracting textual patterns [9], with ESG topic analysis being a key task [54].
+ESG topic analysis facilitates ESG assessments in several ways.
+Specifically, analysing topics allows stakeholders to monitor corporate alignment with UN Sustainability Development Goals [70], to analyse the future orientation of sustainability reports [34], and to construct metrics with respect to ESG performance [68].
+Current NLP methods for this task predominantly include rule-based lexical methods that leverage existing sustainability dictionaries such as Baier [37], and unsupervised topic modeling that can be adapted to ESG analysis [82, 45].
+Yet, while having advantages as automated tools [2], these methods may not be tailored for the specialised requirements of ESG topic analysis.
+Moreover, building specific tools for ESG topic analysis has been under-explored, with efforts primarily focusing around the creation of sustainability lexicons [5, 60, 39].
+However, these lexicons have notable limitations, which we will discuss further in the next section.
+5.2 Lexicons for ESG Analysis
+Lexical approaches have been widely used to analyse sustainability reports [34], given their effectiveness in corporate text analysis and accessibility to stakeholders with diverse technical backgrounds [51, 11].
+To facilitate lexical analysis, lexicons specific to ESG analysis have been developed.
+For instance, approaches such as Baier’s dictionary [5], have focused on developing single-word lexicons according to ESG pillars.
+However, while Baier’s dictionary has been a popular choice for ESG analysis studies [37], its single-word lexicons cannot adequately capture the semantics of the sustainability field [65].
+In contrast, dictionaries that have collected ESG-related text spans (of multiple-grams)—i.e.
+Du et al.
+[23]Kang and El Maarouf [39]—possess limited ESG topics or are not publicly available.
+Additionally, due to their reliance on manual annotation, all the aforementioned dictionaries have an exceedingly small number of lexicons (less than 800).
+This raises questions on their comprehensiveness and consequently their effectiveness for capturing textual patterns.
+To address these limitations, our constructed knowledge base introduces a structured lexicon of ESG concepts—one that includes multiword expressions that more effectively capture the semantics of the sustainability field, and is more extensive, comprising 23k concepts.
+Additionally, existing publicly available ESG lexicons—which, to the best of our knowledge, predominantly include Baier [5] and Naiara [60]—represent structured attempts to derive ESG knowledge, but differ significantly from our neuro-symbolic framework in both methodology and scalability.
+Baier’s dictionary is constructed through manual annotation and term frequency heuristics applied to sustainability documents, limiting its scalability and adaptability.
+Naiara’s dictionary, by contrast, derives ESG-related terms via cosine similarity to a small set of predefined anchors (“environmental,” “social,” and “governance”), which restricts the diversity of retrieved concepts and biases term selection toward those semantically proximal to the anchors.
+In contrast, our approach is fully automated and designed to support both scalability and conceptual diversity.
+It combines GPT-4o-based annotation with label propagation, uses automated concept parsing to extract candidate ESG terms, and applies data clustering and graph-based analysis to select seed terms from semantically diverse regions of the data space.
+Compared to existing lexicons, our framework yields a relatively more diverse and extensive repository of ESG lexicons (or concepts), organised within a structured topical hierarchy.
+As highlighted by our experimental results in the later section (Section 2.3), this provides a greater number of unique ESG terms extracted in ESG topic analysis.
+5.3 ESG Knowledge Graphs
+While prior works have applied knowledge graphs to the analysis of sustainability data [31, 4], these efforts are primarily constructed to support downstream tasks such as relation extraction or question answering over ESG-related news or sustainability reports.
+In these cases, the knowledge graphs are designed to represent document-specific entities and relations, but are not formalized as reusable or publicly accessible repositories of ESG concepts.
+Moreover, these approaches are not oriented toward ESG topic analysis—they do not aim to identify, structure, or surface ESG topics and topical terms systematically across corpora.
+By contrast, ESGSenticNet is explicitly designed as a structured and reusable knowledge base of ESG concepts: it organizes ESG terms and their topical relations into a coherent taxonomy, and is openly available for external use.
+While ESGSenticNet can also support downstream applications such as relation extraction, its foundational role is to serve as a general-purpose, publicly accessible repository for ESG topic analysis.
+Furthermore, because ESGSenticNet can be deployed through a structured lexical method, it requires no training or technical expertise—making it accessible to stakeholders across diverse backgrounds.
+5.4 Semantic Knowledge Bases
+In the literature, semantic knowledge bases are often represented as knowledge graphs consisting of triples in the form of (subject, predicate, object), or (head, relation, tail), which encode structural relations between entities [38].
+By structuring knowledge in a way that mimics human cognition, knowledge bases have been explored in numerous NLP applications, including commonsense question answering [20], sarcasm detection [78], and sentiment analysis [81], among others, as part of broader research efforts toward more context-aware and human-like language understanding.
+However, the construction of knowledge bases within the ESG domain remains largely unexplored.
+To address this gap, this paper presents the first attempt at developing a knowledge base that organises ESG knowledge into structured relationships.
+Methods for constructing knowledge bases can vary according to the domain and application [38].
+Notable knowledge bases, such as ConceptNet [42] involved concept parsing and human annotation, SenticNet [14] leveraged deep learning and lexical substitution, other knowledge bases involve question answering via advanced graphical neural networks (GNNs) [77].
+Newer avenues for knowledge construction involve generative-LLMs, which have shown strong performance in related tasks such as reasoning, construction, and generalisation [83].
+Unlike all the previous methods which heavily rely on semantic similarities for labelling, generative LLMs possess powerful emergent capabilities that extend beyond textual similarities [72].
+These include reasoning, knowledge relationship construction, and generalisation to different domains [83].
+Moreover, unlike GNNs that rely on heavily structured text data [77], generative-LLMs can flexibly adapt to unstructured text [7].
+Yet, leveraging generative LLMs for knowledge construction is relatively new, with limited exploration of its effectiveness.
+This study contributes to this area by constructing a ESG-focused knowledge base using generative LLMs, and evaluating the resulting knowledge base’s accuracy in categorising ESG concepts under the correct topics.
+5.5 Neurosymbolic Methodologies
+Neurosymbolic approaches integrate symbolic AI (rule-based systems) with sub-symbolic AI (neural networks) to enhance generalisation, interpretability, and reasoning in AI [33].
+These methods have been successfully applied to various NLP tasks, including word sense disambiguation for sentiment analysis [79], recommender systems [66], and question-answering [20].
+However, their potential for large-scale knowledge base construction remains largely unexplored.
+This study extends neurosymbolic methodologies by developing a neurosymbolic framework for knowledge construction.
+Specifically, sustainability linguistic patterns are integrated into rule-based concept extraction while GPT-4o is used for labelling concepts’ relations with ESG topics.
+In doing so, this study explores the potential of a neurosymbolic approach to enhance the relevance and accuracy of knowledge construction.
+6 Conclusion
+In this study, we described the challenges of ESG topic analysis, outlining the limitations of existing NLP approaches—i.e., irrelevance, immateriality, limited organization.
+To specifically address these challenges, we developed a knowledge base, ESGSenticNet, from a novel neurosymbolic framework, before evaluating its effectiveness against existing NLP benchmarks.
+The results indicate that ESGSenticNet captures more relevant, material, and comprehensive insights for ESG topic analysis, significantly outperforming existing sustainability dictionaries and state-of-the-art NLP methods.
+Additionally, the topic terms ESGSenticNet provides are correctly associated with their intended topics to a reliable degree, highlighting its robustness for topic analysis.
+All in all, ESGSenticNet’s deployment as a structured lexicon method does not require computational training or technical expertise, meaning it can be readily used by stakeholders regardless of their technical background.
+Ultimately, we hope that ESGSenticNet democratises access to crucial insights for ESG assessments.
+7 Limitations
+Although ESGSenticNet is designed to be an accessible and inclusive tool for ESG topic analysis, it has several limitations.
+First, due to the evolving nature of ESG frameworks [54], ESGSenticNet may require periodic updates to incorporate newly emerging ESG topics.
+Second, the resource is currently limited to concepts expressed in English and is not yet applicable to non-English sustainability reports.
+Future work will therefore focus on extending ESGSenticNet to multiple languages, potentially through the use of multilingual large language models [50].
+Third, while ESGSenticNet covers a broad range of ESG topics and is built on a large corpus of sustainability reports, it does not explicitly assess the generalisability of its concepts across reports from different industries (e.g., finance vs.
+energy).
+In other words, our study does not investigate whether ESGSenticNet provides equally effective topic analysis for sustainability reports from companies in different sectors.
+Addressing this may involve developing cross-industry generalisation tasks and exploring learning algorithms (i.e.
+contrastive learning) to improve robustness [55]
+Additionally, while our neurosymbolic approach mitigates overreliance on LLMs (specifically GPT-4o) by incorporating sustainability-specific linguistic patterns through symbolic concept parsing, we acknowledge that LLMs may still exhibit category-specific biases due to their pretraining [21].
+For instance, they may be more accurate in classifying relations for certain sustainability categories over others [55].
+To address this, we decompose the LLM inference task into smaller, focused sub-tasks (Section 4.3.3), improving accuracy even with less familiar categories [17].
+We further validate the reliability of LLM classifications through a human evaluation study detailed in Section 5.4.
+Declarations
+The authors declare that they have no competing interests as defined by Nature Portfolio, or other interests that might be perceived to influence the results and/or discussion reported in this paper.
+Author Contributions: K.
+O.
+: Writing – review & editing, Writing – original draft, Methodology, Investigation, Formal analysis, Conceptualisation.
+R, M.: Writing – review & editing, Methodology, Investigation, Concep- tualisation.
+D.V.: Conceptuslisation, F.X: Conceptualisation.
+R.
+S.: Resources.
+J.
+S.: Writing – review & editing, Methodology, Investigation.
+E.
+C.
+Writing – review & edit- ing, Conceptualisation, Supervision.
+G.
+M.: Writing – review & editing, Supervision, Methodology, Conceptualisation.
+Acknowledgements & Funding Statement
+This research/project is supported by the NUS Sustainable and Green Finance Institute (SGFIN), NUS Asian Institute of Digital Finance (AIDF), Ministry of Education, Singapore under its MOE Academic Research Fund Tier 2 (STEM RIE2025 Award MOE-T2EP20123-0005), MOE Tier 2 Award (MOE-T2EP50221-0006: “Prediction-to-Mitigation with Digital Twins of the Earth’s Weather”), MOE Tier 1 Award (MOE-T2EP50221-0028: “Discipline-Informed Neural Networks for In- terpretable Time-Series Discovery”), and by the RIE2025 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) (Award I2301E0026), administered by A*STAR, as well as supported by Alibaba Group and NTU Singapore.
+Appendix A ESGSenticNet Taxonomy
+Table 16 contains the definitions of the pillars (Environmental, Social, Governance) and topics used in ESGSenticNet.
+Along with the pillars, the broad topics, and cross-broad topics are bolded, with their constituent sub-topics highlighted in bullet points.
+Beside the broad and cross-broad topic names, a letter denotes which pillar the topic falls under – (E) Environmental, (S) Social, (G) Governance.
+Topic Description
+Environmental Focuses on impact on the environment.
+This includes, but is not limited to, green innovation, green infrastructure, resource stewardship, energy consumption, waste management, pollution reduction, natural resource conservation, and animal welfare.
+Social Focuses on managing relationships with its employees, suppliers, customers, and the communities.
+This includes, but is not limited to, employee diversity, equity, working conditions, customer relations, partnerships and social engagement.
+Governance Focuses on the system of rules, practices, and processes by which a company is directed and controlled.
+This includes, but is not limited to, corporate oversight, corporate transparency, executive compensation, ethical behavior, and shareholder rights.
+Development (E) Drives innovation and growth in sustainability by creating new technologies or infrastructure for sustainability, or supporting these efforts through green financing.
+$\bullet$ Green Technologies Technological systems, products, and processes that assist in reducing environmental impact.
+$\bullet$ Green Infrastructure Environmentally sustainable buildings, infrastructure, energy solutions and urban systems.
+$\bullet$ Green Financing Financial services that support efforts aimed at environmental sustainability.
+Ecological Conservation (E) Manages the company’s interaction with natural ecosystems, focusing on minimizing harmful impacts and promoting biodiversity and habitat conservation.
+Table 12: ESGSenticNet Sustainability Taxonomy - continued in the next page
+Emissions Control (E) Targets the reduction of various emissions from company operations, with an emphasis on mitigating climate change impacts or protecting air quality.
+$\bullet$ Climate Emissions Manages greenhouse gas emissions that contribute to global warming, such as CO2 and methane.
+The phrase must explicitly relate to greenhouse gases or related terms such as ’carbon’, ’methane’.
+$\bullet$ Air Quality Emissions Manages emissions of non-greenhouse gases, like toxic gas, particulates and sulfur dioxide, which are critical for local air quality and public health.
+Operations (E) Focuses on optimizing existing business activities for the sole purpose of minimizing environmental impact.
+This includes minimising environmental impact by enhancing current practices in corporate operations, supply chains, production processes, or land management, distinct from the development of innovative solutions and technologies.
+$\bullet$ Sustainable Corporate Operations Environmentally responsible practices within corporate offices.
+Examples include but are not limited to promoting sustainability amongst employees, eco-friendly office lighting.
+$\bullet$ Sustainable Supply Chain Logistics and distribution methods or practices that are environmentally responsible, ensuring sustainability in the movement and handling of goods.
+$\bullet$ Sustainable Production Processes Manufacturing processes or practices that are environmentally responsible.
+$\bullet$ Sustainable Land Management Using and managing land in a way that is environmentally responsible, conserves natural resources and protects biodiversity.
+Waste Management (E) Concentrates on managing waste after it has been produced, including the reduction, handling, disposal, and treatment of solid waste and wastewater.
+It involves but is not limited to recycling, reusing materials, and employing wastewater treatment processes to mitigate environmental degradation and facilitate resource recovery.
+$\bullet$ Wastewater Management Focuses on the treatment, reuse and recycling of wastewater or liquid waste to prevent pollution or conserve water resources.
+Table 13: ESGSenticNet Sustainability Taxonomy - continued in the next page
+$\bullet$ Solid Waste Management Addresses the management of solid wastes, promoting reduction, recycling treatment, and responsible disposal to minimize environmental impact and encourage circular economy principles.
+Resource Optimisation (E) Prioritizes the proactive, efficient, and sustainable selection and use of energy, water, and materials.
+This approach focuses on minimizing environmental impact through the strategic choice and utilization of resources.
+Distinct from waste management as it does not primarily deal with waste but rather prevents waste generation by optimizing resource use from the start.
+$\bullet$ Renewable and Efficient Energy Focuses on using energy efficiently, reducing energy consumption and increasing the use of renewable energy sources such as solar, wind, and hydroelectric power to minimize environmental impact.
+$\bullet$ Material Sustainability Prioritizes the use of sustainable and renewable materials, focusing on maximizing their efficiency from the outset.
+Distinct from recycling and reuse, which are part of waste management, this approach aims to reduce environmental impact before materials enter the waste stream.
+$\bullet$ Water Conservation Focuses on using water efficienctly, reducing water usage and enhancing sustainable water management practices to minimize initial consumption.
+Distinct from water recycling and reusing, which address management of wastewater.
+Workplace (S) Focuses on the social aspects of sustainability within the company to promote equity, inclusivity, or quality of life, emphasizing employee well-being, inclusivity, or professional growth.
+$\bullet$ Workplace Wellness Efforts that focus on the satisfaction, emotional and physical well-being of employees.
+$\bullet$ Workplace Diversity Efforts that emphasise gender and racial diversity and fairness in the workplace, such as anti-discrimination measures, fair hiring practices etc.
+$\bullet$ Workplace Development Efforts that focus on the skill training and professional development of staff.
+Table 14: ESGSenticNet Sustainability Taxonomy - continued in the next page
+Communications (E/S/G) Focuses on how a company transparently and honestly communicates and discloses its activities.
+Outreach (S) Enhances social sustainability by contributing to community well-being, fostering customer satisfaction, building customer relationships, or collaborating with strategic partners to promote equity, inclusivity, and quality of life.
+$\bullet$ Community Empowerment Contributions to the welfare of local communities through various forms of support and engagement.
+$\bullet$ Strategic Partnerships Collaborations with stakeholders like NGOs, businesses, and charities to achieve shared sustainability goals.
+$\bullet$ Customer Engagement Fostering customer satisfaction, strong relationships and reputation with customers.
+Management (G) Ensures strategic, ethical, or effective governance through rigorous oversight, diverse board composition, active stakeholder engagement, or high ethical standards within company leadership.
+$\bullet$ Management Composition Structures the board or management with diverse expertise and backgrounds to strengthen decision-making and governance effectiveness.
+$\bullet$ Management Ethics Enforces strict ethical standards within company leadership, ensuring management actions uphold corporate integrity.
+$\bullet$ Management Compensation Aligns management, executive and board remuneration with company performance, sustainability or ethical objectives.
+$\bullet$ Stakeholder Engagement Engages key stakeholders to align board or management decisions with broader interests and feedback.
+$\bullet$ Oversight Involves the systematic review and evaluation of management actions to ensure they align with the company’s performance goals, sustainability goals and ethical standards.
+Compliance (G) Ensures adherence to laws, regulations, or standards across environmental protection, labor rights, consumer safety, ethical conduct, financial practices, or data protection, distinguishing itself from voluntary sustainability initiatives by focusing on mandatory requirements.
+Table 15: ESGSenticNet Sustainability Taxonomy - continued in the next page
+$\bullet$ Environmental Compliance Adherence to environmental laws such as for emissions, water pollution, protection of biodiversity etc.
+$\bullet$ Worker & Consumer Safety Adherence to laws that guarantee the safety of products and services to consumer health, as well as fair labor practices and the protection of workers’ rights.
+$\bullet$ Business Compliance Adherence to anti-corruption, anti-competitive practices, financial regulations, or financial reporting.
+$\bullet$ Data Privacy & Cybersecurity Protection Protection and confidentiality of data, which includes measures against cyber threats and compliance with data security regulations.
+Table 16: ESGSenticNet Sustainability Taxonomy
+Appendix B Additional details on Human Validation
+For evaluating ESGSenticNet accuracy (Section 2.4), ESGSenticNet’s flexible matching method (Section 4.10.3), GPT-4 human agreement (Section 4.10.2), there are a total of 9 human annotators.
+These annotators are based in Singapore, comprising researchers from the Asian Institute of Digital Finance and the Sustainable and Green Finance Institute, who are pursuing doctoral-level and post-doctoral research within the corporate sustainability field.
+All human annotators contribute to this study as part of their formal academic and research responsibilities.
+All human validation studies follow the same annotation scheme.
+When there are disagreements, discussions are held to achieve a resolution.
+Where there is no clear resolution, majority voting is taken to determine the ground truth label.
+For the derivation of the taxonomy for ESGSenticNet, we involved two tenured researchers from the Singapore Sustainable and Green Finance Institute (SGFIN), with expertise within the sustainable finance field.
+Appendix C Full Statistical Breakdown of ESGSenticNet
+Topics Total triples Number of triples with the relations
 supports undermines aligns with
@@
 Table 17: ESGSenticNet Statistics
-Concept Relation Category
-build diverse workplace supports workplace diversity
-halve carbon emission supports emissions control
-involve workplace injury undermines worker & consumer safety compliance
-minimise resource consumption supports resource optimisation
-organise charity event supports outreach
-produce chemical waste undermines operations
-Table 7: Samples of knowledge triplets from ESGSenticNet
-Samples of knowledge triplets within ESGSenticNet is provided in table 7. Additionally, a full statistical breakdown of ESGSenticNet’s 44k knowledge triplets is outlined in table 6.
-5.9 ESGSenticNet Human Evaluation
-Sustainability Corpora Total Terms
-ESGSenticNet 23245
-Baier 491
-Naiara 167
-Kang* 792
-Table 8: Total number of terms of different sustainability corpora, Baier [6], Naiara [44], Kang [29], *indicates not publicly available.
-Relation Type Accuracy
-Pillar 95.0%
-Broad 89.3%
-Cross-Broad 86.1%
-Sub 87.8%
-Cross-Sub 90.1%
-Table 9: Accuracy of ESGSenticNet
-ESGSenticNet comprises 44232 triplets containing 23245 unique concepts.
-Of the 23245 concepts, 16011 are seeds and 7234 are non-seeds.
-We provide a meta statistic of ESGSenticNet (table 8), a full statistical breakdown (table 6), and samples (table 7).
-To evaluate our knowledge base, human annotators determine the accuracy of relations between concept and categories.
-500 concepts are randomly selected from our knowledge base.
-A concept may be found in multiple knowledge triplets, possessing multiple relations with different categories.
-The accuracy of the labels are human evaluated at every GPT-4o decision point as shown in figure 4.
-Specifically, we evaluate the accuracy of relation labels separately for each category type–pillar, broad, cross-broad, sub, cross-sub, with each of them including only select categories (table 4).
-A provides more details on our annotators and annotation scheme.
-For evaluating ESGSenticNet accuracy in section 5.9, the following instructions were given:
-The table below presents various concepts along with their polarities and respective relations toward sustainability categories.
-Each concept may be associated with multiple relations across different sustainability categories.
-These categories are classified into five category types: pillar, broad, sub, cross-broad, and cross-sub.
-If a concept has no relevant association with a category, it is marked as ‘not applicable’.
-Your task is to evaluate whether each concept is correctly related to its corresponding categories.
-For each category, determine if the relation is the most appropriate compared to the other categories within the same category type.
-Mark ‘TRUE’ if the relation and category are correct, and ‘FALSE’ if they are incorrect.
-Additionally, assess the accuracy of the polarity of each concept, and mark ‘TRUE’ if it is correct and ‘FALSE’ if it is incorrect.
-6 Topical Analysis of Sustainability Disclosures through ESGSenticNet
-Topic analysis of sustainability disclosures provide stakeholders significant insights about a corporation’s sustainability performance [41].
-Distinct from discovering topics without any inclination of what the topics should be, topic analysis within sustainability analysis frequently seeks to derive topics within established sustainability themes and frameworks [41].
-To elaborate on this, sustainability stakeholders often have specific themes they are monitoring, such as alignment with UN Sustainability Development Goals [57], future orientation of reporting within ESG pillars [23], performance with respect to economic, environmental, social dimensions [54].
-Therefore, the demands of the topic analysis within sustainability requires capturing content related to sustainability and its associated frameworks.
-ESGSenticNet aims to address the shortcomings with respect to unstructured topic terms, unstructured topics, the limited structured methods, as well as the materiality of information extracted.
-To elaborate, the hierarchical structure of ESGSenticNet ensures that the topics extracted are directly related to a sustainability taxonomy, while the specialised concept parsing and GPT-4o inference increases the materiality and ESG-relevance of extracted topic terms.
-6.1 Critique of current methods
-Different methods can be leveraged for sustainability topic analysis, including rule-based methods that leverage existing dictionaries [25], and unsupervised topic modelling that can be adapted to explore specific areas within sustainability frameworks [66, 35].
-Yet, while having advantages, these methods may not be tailored for sustainability topic analysis, given the specialised nature of sustainability content and the demands of sustainability stakeholders.
-We elaborate on this in the following.
-  • Unrelated topic terms: Existing topic modelling methods, while excellent at producing topic terms without supervision  [1], cannot be tuned to capture specialised word constructions that typically describe sustainability ideas (i.e.
-nominals, modifiers as described in 5.2[50].
-Moreover, sustainability content is multifaceted and complex, often varying across different sectors and contexts [39].
-Therefore, captured topic terms may not be relevant to sustainability, leading to topics irrelevant for sustainability analysis.
-  • Unstructured topics: Unsupervised topic modelling may not yield topics that are structured according to sustainability themes, even if sustainability-related topics can be derived from topic words.
-Therefore, the derived topics may not be meaningful for sustainability stakeholders, who typically rely on structured frameworks with predefined sustainability themes and topics to analyse sustainability content [34, 19].
-  • Limited Structured Methods: Although publicly available ESG dictionaries such as Baier [6] introduce a structured method for topic analysis [22], they contain lexicons, which we will demonstrate, that are not predominantly related to sustainability.
-This makes the alignment between the lexicons and their supposed sustainability categories unclear.
-Other dictionaries (i.e Naira [44]) are limited ($<200$ terms).
-  • Material Information: Although topic terms may highlight the presence of sustainability issues, stakeholders often prioritise content that explicitly convey the sustainability efforts or actions undertaken by corporations [7].
-Therefore, greater specificity of topic terms is required to enrich the materiality of topical information in alignment with stakeholder interests.
-Nevertheless, current methods may fall short in generating specialised topic terms that articulate sustainability action, a limitation that we will explore further.
-6.2 Topic Analysis Evaluation
-In line with the aforementioned critiques and requirements with respect to sustainability topic analysis, ESGSenticNet is tested against baselines.
-These include, topic models that are algebraic – NMF [30], probabilistic – LDA [9], neural – ProdLDA [52], TSCTM [60], BERTopic [21], ECRTM [59], and other publicly available sustainability dictionaries – Baier [6] Naiara [44].
-We aim to answer the following research questions to assess whether the tested methods provide meaningful topic analysis for sustainability stakeholders:
-  • RQ1: To what extent are the topic terms yielded by a method related to ESG, and by extension, sustainability analysis?
-  • RQ2: To what extent are the ESG topic terms yielded by a method diverse, allowing analysts to view a variety of sustainability topics?
-  • RQ3: To what extent do the topic terms yielded by a method express an action toward ESG, or a sustainability effort, thereby providing more material insights?
-To address these research questions, topic analysis methods are evaluated according to the following.
-  • ESG relatedness: The proportion of all topic terms that are ESG-related, highlighting the extent to which yielded topics are relevant to sustainability and therefore sustainability analysis.
-  • Unique ESG terms: The number of unique ESG topic terms, indicating the comprehensiveness and diversity of ESG content captured.
-  • ESG action orientation: The proportion of all topic terms that express an action taken toward ESG, or in other words, a sustainability effort.
-This indicates whether the terms are able to express information material, and of interest to stakeholders.
-For the above metrics, GPT-4 is leveraged to classify all topic terms yielded from the models.
-GPT-4 distinguishes the following – 1) whether each topic term is ESG-related, 2) whether each topic term constitutes an action toward improving ESG.
-Thereafter, we aggregate the GPT-4 classification results to produce scores for ESG relation, unique ESG terms, ESG action.
-We validate using GPT-4 in our evaluation through an GPT4-human agreement study (section 6.6).
-Full experimental details for topic analysis evaluation are described in section 6.3, with the prompts provided in figure 7 (ESG relatedness), and figure 8 (ESG action orientation).
-[Task Description] As a sustainability expert, your task is to classify terms based on their relevance to Environmental, Social, and Governance (ESG) factors.
-[Response Instructions] Evaluate each term based on its content to determine its most appropriate category.
-If a term is explicitly related to ESG, assign the label ‘ESG’.
-If a term is not related to ESG or the relationship is vague, assign the label ‘non-ESG’.
-Ensure all responses strictly adhere to the definitions and format provided.
-[Response Format] Fill in your response in the triple backticks below using the following format: Input: <term> Response: <ESG or non-ESG>
-Figure 7: Prompt for evaluating ESG relatedness
-[Task Description] As a sustainability expert, your task is to classify terms based on whether they express an action taken toward ESG.
-[Response Instructions] Evaluate each term based on its content to determine whether it describes an action taken to advance Environmental, Social, and Governance (ESG) goals.
-If a term describes an action taken to enhance ESG objectives, assign the label ‘True’.
-If a term does not describe an action taken toward ESG objectives, assign the label ‘False’.
-Ensure all responses strictly adhere to the definitions and format provided.
-[Response Format] Fill in your response in the triple backticks below using the following format: Input: <term> Response: <True or False>
-Figure 8: Prompt for evaluating ESG action orientation
-6.3 Experimental Settings
-Topic analysis is performed on 319 sustainability disclosures from 75 large Singaporean Companies from 2015 to 2023, obtained from the respective company websites.
-These reports were unused during the development of ESGSenticNet.
-Algebraic, Probabilistic & Neural methods such as NMF [30], LDA [9], ProdLDA [52], BERTopic [21], TSCTM [60], ECRTM [59], are unsupervised, and we run them across a wide range of hyperparameters, to obtain the best results for each method.
-These parameters include every combination for vocab size=$=\{2500,5000,7500,10000,12500\}$, number of topics=$=\{10,20,30,40,50,60,70\}$, number of topic words=$=\{5,10,15,20,25,30\}$.
-Additionally, these models were deployed through the TopMost package [61].
-Dictionary methods comprise terms that are categorised according to specific sustainability topics.
-A topic’s occurrence is computed from the frequency of occurrence of its associated terms.
-In our experiment, the topic terms of the dictionaries correspond to dictionary terms that appear within the corpora of sustainability disclosures.
-To the best of our knowledge, dictionary methods in the published literature are limited, with only a few works that make their full dictionaries publicly available.
-These include Baier [6] and Naiara [44].
-ESGSenticNet comprises knowledge triplets in the format (concept, relation, category), delineating how specific concepts related with sustainability categories or topics.
-To enhance semantic specificity and context, we employ three word concepts, exclusively using those that possess the ‘supports’ relation toward a sustainability topic.
-Similar to dictionary methods, a topic’s occurrence is determined by the frequency of appearance of concepts that support the topic (we explore this further in a case study in section (6.5)).
-The topic terms of ESGSenticNet correspond to these concepts that appear within the sustainability disclosures corpora.
-Additionally, we experiment with two configurations for ESGSenticNet: Exact involves exact string matches of each concept, Flexible involves taking advantage of the (verb, noun phrase) arrangement of concepts for matching.
-To elaborate on this, the verb (1st word) is matched separately from the corresponding noun-phrase (2nd & 3rd words) within each sentence.
-Given the example “We reduce our water consumption”, ‘reduce’ is matched as the verb, and ‘water consumption’ is matched as the noun-phrase, with ‘reduce water consumption’ returned as the matched concept without requiring its exact string match.
-Flexible enables greater detection of ESG concepts by allowing for semantic variability (see study in section 6.7), while Exact is useful for detecting precise ESG impacts.
-6.4 Results & Discussion
-Baseline Best ESG-unique Best ESG-rel Best ESG-act
-ESG-unique ESG-rel ESG-act ESG-unique ESG-rel ESG-act ESG-unique ESG-rel ESG-act
-NMF 332* 0.28* 0.08* 3 0.42 0.00 332* 0.28* 0.08*
-LDA 23* 0.31* 0.01* 3 0.36 0.00 23* 0.31* 0.01*
-ProdLDA 334* 0.26* 0.07* 17 0.38 0.10 13 0.38 0.16
-BERTopic 76 0.33 0.05 13 0.40 0.04 13 0.35 0.05
-TSCTM 359 0.15 0.06 11 0.32 0.06 23 0.31 0.15
-ECRTM 440 0.18 0.09 23 0.46 0.08 182 0.31 0.15
-Baier 212* 0.49* 0.29* 212* 0.49* 0.29* 212* 0.49* 0.29*
-Naiara 1* 0.50* 0.50* 1* 0.50* 0.50* 1* 0.50* 0.50*
-ESGSenticNet (exact) 359* 0.76* 0.84* 359* 0.76* 0.84* 359* 0.76* 0.84*
-ESGSenticNet (flexible) 2555* 0.79* 0.81* 2555* 0.79* 0.81* 2555* 0.79* 0.81*
-Table 10: Evaluation of Topic Analysis Methods, through the metrics of unique ESG terms (ESG-unique), ESG relatedness (ESG-rel), and ESG action orientation (ESG-act).
-Best results are marked in bold and the best baseline results are underlined.
-(*) indicates the results from the same run (i.e.
-same parameters).
-Best ESG-unique, Best ESG-rel, and Best ESG-act represent the top-performing results from the respective runs of the tested methods.
-Our experimental results (table 10) indicate that ESGSenticNet scores higher than the other topic analysis methods on ESG relatedness and ESG action orientation by at least 26% and 31% respectively.
-Additionally, both ESGSenticNet (exact) and ESGSenticNet (flexible) extract a high number of unique ESG terms, with ESGSenticNet (flexible) scoring the highest amongst all methods tested.
-Although requiring limited computational resources as a lexical method, ESGSenticNet produces significantly higher performance compared to the more expensive unsupervised topic modelling methods.
-This is despite how the topic models were exhaustively ran on numerous permutations of parameters (section 6.3) that required extensive computational costs.
-Unsupervised topic models such as TSCTM, ECRTM, ProdLDA, NMF can extract a high number of unique ESG terms ($>300$).
-Yet, for the same runs, the concentration or prevalence of ESG terms within each topic is modest, as highlighted by their limited scores in ESG relatedness ($<0.30$).
-Therefore, despite capturing diverse ESG topic terms, these methods fall short in generating topics that are richly centered on ESG themes.
-From table 11, we highlight examples of the topics generated by these methods in comparison to ESGSenticNet, highlighting how most of the topics generated by these topic models are not ESG-centric.
-This limitation is critical in sustainability topic analysis, given that the primary goal is to derive insights from topics that are not only inclusive of ESG terms but are also substantially focused on ESG issues.
-Moreover, topic model runs that yield the highest ESG relatedness are only able to yield few unique ESG-terms ($<25$).
-Furthermore, for these topic model configurations, as well as for other dictionaries such as Baier and Naiara, ESG relatedness is still $\leq 0.50$.
-This suggests that yielded topics are still not predominantly centred around ESG themes.
-Finally, most of the tested baseline methods score modestly for ESG action orientation, Naiara noticeably higher.
-Yet, Naiara is only able to derive $1$ unique ESG-term, which limits the significance of its analysis.
-These results suggest that the baseline methods may not fully capture information that is material to stakeholders, particularly in terms of conveying sustainability actions and efforts.
-Method Topic Example Topic Words
-TSCTM Topic 1 offshore, baker, vessels, oil, cho, vessel, chartering, ship, sea charter
-Topic 2 rsp, venture, johor, black, sites
-Topic 3 controlling, provision, shall, period, accounting, mainly, listing, media, entity, time
-ECRTM Topic 1 minerals, healthcare, disa, teh, bhd holdings, agm, sdn, heatec, sri
-Topic 2 banks, grand employees, duty, palm
-Topic 3 disa, shares, theft share, anniversary, lau, purchase, options, june cancellation
-ProdLDA Topic 1 financial, value, non assets, december
-Topic 2 loss, company, board, statements, directors
-Topic 3 greenhouse, mountain, aspires, gresb, honeyness, feel, strategy, board, candid, phenomenon, conforms, performance, cubic, cabling, unoccupied
-NMF Topic 1 duly, assets, transparency, value, priority
-Topic 2 property, fair, rate, million, new
-Topic 3 esg, administration, confirmed, quarter, save, present, businesses, employees, workplace, flow, focusing, ceo, default, meetings, termination
-ESGSenticNet Topic 1 abide local law, achieve regulatory compliance, fulfill regulatory requirement, follow international law, comply environmental rule
-Topic 2 increase employee satisfaction, improve staff engagement, invest career development, maintain employee safety
-Topic 3 minimise carbon footprint, monitor ghg emission, reduce environmental pollution, lower air emission
-Table 11: Examples of topic words for topic models with the highest numbers of unique ESG terms, in comparison with ESGSenticNet for qualitative analysis
-6.5 Case Study - Insights offered by ESGSenticNet
-Figure 9: Frequency of ESGSenticNet topics from the flexible run
-Figure 10: Top ESGSenticNet concepts from the flexible run, in support of the sub topics (dark blue), and their corresponding broad topic (light blue)
-The topic analysis conducted by ESGSenticNet in section 6 is visualised with the figures 9 & 10, which represent different ways to analyse the text corpus.
-Figure 9 provides a higher level breakdown of the topics present in the corpus.
-It highlights the top 5 frequently occurring broad sustainability topics present in the corpus, as well as the proportion of their constituent sub topics, in accordance with ESGSenticNet’s hierarchical taxonomy.
-The topic frequencies are denoted by the occurrence of matched concepts that ‘support’ a topic, leveraging ESGSenticNet’s predefined relations.
-Separately, figure 10 provides a more granular analysis of the corpus, highlighting matched concepts that explicitly convey sustainability actions, thereby communicating important sustainability information.
-Figure 10 displays the top occurring concepts that support the topics shown in figure 9.
-Put together, figures 9 and 10 highlight different methods to analyse a sustainability corpus, which can be utilised depending on a stakeholder’s interest and preference.
-6.6 GPT-4 Human Expert Agreement on Topic Analysis Evaluation Metrics
-GPT-4 evaluations have shown significant reliability in the literature [64].
-A study is conducted to assess GPT-4’s ability for the following tasks– i) classifying topic terms as ESG-related, ii) classifying topic terms as actions taken toward ESG.
-A random sample of 510 topic terms were gathered for each task.
-GPT-4 annotations for each set of topic terms are compared with annotations from 2 different groups of 3 human experts each, with each group handling a different task.
-An average agreement score of 83.7% observed for task i) and 91.8% for task ii).
-For evaluating GPT-4 human agreement, the following instructions were given for classifying topic terms as ESG-related.
-The table below presents various terms. Determine if each term is ESG-related. Mark ‘TRUE’ if the term is ESG-related and ‘FALSE’ otherwise.
-For evaluating GPT-4 human agreement, the following instructions were given for classifying topic terms as an action taken toward ESG.
-The table below presents various terms.
-Determine if each term expresses an action toward improving a company’s ESG performance.
-Mark ‘TRUE’ if the term express an action toward improving ESG performance and ‘FALSE’ otherwise.
-6.7 Effectiveness of ESGSenticNet Flexible Matching Method
-A study is conducted to assess the precision of matching ESGSenticNet concepts through the ‘flexible’ method.
-A random sample of 200 sentences matched through this method was collected, excluding sentences matched via the ‘precise’ method.
-Human annotators were tasked with evaluating whether each sentence clearly conveys the idea expressed represented by its matched concept.
-From our results, 75.0% of sentences comprised the idea expressed by its corresponding matched concepts.
-This suggests that the ‘flexible’ method retains a moderately high level of precision, despite allowing for greater semantic variability.
-For evaluating the ESGSenticNet flexible matching method, the following instructions were given to evaluate if each sentence clearly conveys the idea expressed by its matched concept.
-The table below presents different sentences and a corresponding concept.
-Determine if each sentence clearly conveys the idea expressed by its matched concept.
-Mark ‘TRUE’ if the sentence clearly conveys the concept’s idea, and ‘FALSE’ otherwise.
-7 Conclusion & Future Work
-This paper suggests that by focusing on the identified challenges of immateriality, complexity and subjectivity, NLP methods can be enhanced for sustainability analysis.
-As our findings indicate, ESGSenticNet, developed with these challenges in mind, provides more relevant and actionable insights for sustainability stakeholders, presentiing a significant improvement over existing sustainability lexicons, let alone state-of-the-art NLP methods for corporate sustainability analysis.
-This is despite how ESGSenticNet’s deployment does not entail computational training or technical expertise.
-Ultimately, ESGSenticNet aims to be a tool that can be readily used by stakeholders regardless of their technical background, to democratise access to crucial corporate sustainability insights.
-Appendix A Additional details on Human Validation
-For evaluating ESGSenticNet accuracy (section 5.9), ESGSenticNet’s flexible matching method (section 6.7), GPT-4 human agreement (section 6.6), there are a total of 9 volunteer human annotators.
-These annotators are based in Singapore, comprising researchers from the Asian Institute of Digital Finance, Sustainable and Green Finance Institute, who are pursuing doctoral-level and post-doctoral research within the corporate sustainability field.
-All human validation studies follow the same annotation scheme.
-When there are disagreements, discussions are held to achieve a resolution.
-Where there is no clear resolution, majority voting is taken to determine the ground truth label.
-For the derivation of the taxonomy for ESGSenticNet, we enlisted the help of two tenured researchers from the Singapore Sustainable and Green Finance Institute (SGFIN), whom we deem as experts within the sustainable finance field.